### v2

# FIT-GNN: Faster Inference Time for GNNs Using Coarsening

Abstract
Scalability of Graph Neural Networks (GNNs) remains a significant challenge, particularly when dealing with large-scale graphs. To tackle this, coarsening-based methods are used to reduce the graph into a smaller graph, resulting in faster computation. Nonetheless, prior research has not adequately addressed the computational costs during the inference phase. This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during both training and inference phases. We demonstrate two different methods (Extra-Nodes and Cluster-Nodes). Our study also proposes a unique application of the coarsening algorithm for graph-level tasks, including graph classification and graph regression, which have not yet been explored. We conduct extensive experiments on multiple benchmark datasets in the order of $100K$ nodes to evaluate the performance of our approach. The results demonstrate that our method achieves competitive performance in tasks involving classification and regression on nodes and graphs, compared to traditional GNNs, while having single-node inference times that are orders of magnitude faster. Furthermore, our approach significantly reduces memory consumption, allowing training and inference on low-resource devices where traditional methods struggle.
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work
[…]
images/img1
Figure 1: The figure shows how the coarsening algorithm reduces a graph $G$ with $15$ nodes to a coarsened graph $G^{\prime}$ with $3$ nodes. It also shows which nodes belong to which partition.
1 Introduction
Graph Neural Networks (GNNs) have proven to be a tool of surprising versatility and modeling capacity. However, there are still a few issues that limit their general applicability, the chief among these being the scalability of the GNN models. This scalability issue is problematic for both the training and the inference phases. Previous research has attempted to address this by reducing the graph into a smaller graph in order to reduce the computational cost. One such broad technique is graph coarsening methods, which judiciously remove nodes and edges from the graph. However, there are still certain points of weakness of these algorithms. For one, coarsening methods can potentially remove important node information. Secondly, and more pertinent to the current discussion, coarsening methods only reduce the training computation; they do not address the inference time. The next obvious question would be to ask why we cannot use coarsening methods for inference as well. Given a graph $G$, coarsening algorithms create a coarsened graph $G^{\prime}$. While $G^{\prime}$ is smaller than $G$, the nodes in $G^{\prime}$ are combinations of nodes in $G$, which can be both train and test nodes. As a result, the prediction per test node is not available. In another scenario, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes. Also, let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively. The coarsening algorithm will take the majority label and assign class 00 to $v^{\prime}$. Hence, the model will be trained to predict only the class 00, leading to discarding the quantification of the model’s performance on predicting less represented nodes. This is the research gap that we seek to address in this work. We use coarsening algorithms to partition the graph into a set of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. We show methods to train on these subgraphs that show equivalent results to non-coarsened baselines and sometimes perform better. This comes hand in hand with a significant reduction in inference time and memory consumption.
We demonstrate two different methods to append additional nodes in the subgraphs— Extra Nodes and Cluster Nodes that respectively append additional nodes to each subgraph, either by adding neighboring nodes, or by adding representational nodes corresponding a neighboring cluster. We theoretically characterize the upper bound of the coarsening ratio for which the time complexity of our approach is better than the naive uncoarsened approach. In node classification and node regression tasks, the baselines typically require the entire dataset during inference to get a prediction for a single node. In contrast, our method only needs to pass the subgraph to which the node belongs. Our results show up to $\mathbf{100\times}$ reduction in inference time and up to $\mathbf{70\times}$ less memory than the baseline while achieving equivalent or better performance.
2 Preliminaries
2.1 Graph Notations
Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix. The feature of each node $v_{i}$ in the graph is the $i$th row vector in $X$ represented as $x_{i}\in\mathbb{R}^{d}$. Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges. Let us assume $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$. $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$. We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
2.2 Graph Neural Network
Graph Neural Network is a neural network designed to work for graph-structured data. The representation of each node in the graph is updated recursively by aggregation and transforming node representations of its neighbors. Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
$X_{l+1}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X_{l}\mathcal{W}_{l})$ (1)
where $X_{l}$ is the node representation after $l$ layers of GCN, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is an identity matrix of same dimension as $A$. $\mathcal{W}_{l}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
2.3 Graph Coarsening and Graph Partitioning
Loukas (2018) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$. We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes. Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$. We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes. Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$.
For the coarsened graph, the adjacency matrix $A^{\prime}=P^{T}AP$. Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$. Huang et al. (2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{j,j}=|C_{j}|,j=1,2,\dots,k$.
Along with a coarsened graph, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$. The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$. Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$. $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
3 Related Work
The goal of reducing computation costs has been tackled via different methodologies. Research like Chiang et al. (2019) and Zeng et al. (2020) use a subgraph sampling-based approach. They sample a small subgraph at each training iteration and train the GNN. Other techniques like layer-wise sampling along with mini-batch training have been studied by  Chen et al. (2018); Chen and Zhu (2018); Cong et al. (2020); Zou et al. (2019). Fahrbach et al. (2020) introduced a new coarsening algorithm based on Schur complements to generate embeddings for vertices in a large graph. Huang et al. (2021) introduced using GNN for a coarsened graph. The motivation was to utilize graph coarsening algorithms to reduce the number of nodes and edges in the graph, resulting in less training time. This approach helps to train GNN for graphs with a large number of nodes. However, inference time is not reduced because the nodes cannot be removed during inference. This training process also caused the removal of important node-level information. Later, Xue et al. (2022) explored shifting the training process from graph-level training to subgraph-level. Coarsening algorithms were used to split the graph into $k-$subgraphs, resulting in the removal of a few edges. Further, to speed up the training process, they propose a multi-device training approach where subgraphs are sent to different devices. This results in a faster training process. Kumar et al. (2023) further proposed the usage of feature-based coarsening algorithms.
4 Methodology
In Huang et al. (2021), a graph $G=(V,E,X,W)$ is reduced to $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ using a partition matrix $P$ generated from the coarsening algorithm. Here $X^{\prime}=P^{T}X$. The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task. While using $\arg\max$, some essential information is lost. Therefore, we follow a subgraph-level training process where we don’t remove these label information. To explore another node-level task, such as node regression, functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets. Hence, our goal is to use subgraphs for training and inference purposes.
To tackle the issue of loss of information corresponding to the nodes (and their edges) present at the periphery of the subgraphs after partitioning the graph, we append nodes in two ways:
[…]
images/img2
Figure 2: This figure shows the methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
$\bullet$ Extra Nodes: Adding the $1$-hop neighbouring nodes in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$
However, we don’t perform backpropagation of the prediction on $\mathcal{E}_{G_{i}}$. Xue et al. (2022) introduced this approach to reduce information loss after partitioning. An edge between two Extra Nodes in $\mathcal{E}_{G_{i}}$ is added if they were connected in $G$.
Algorithm 1 Node-Model$(L,A,D,X)$
1:Number of Layers $L$; $A$; $D$; $X_{0}=X$;
2:for $i=1$ to $L$ do
[…]
4:end for
5:$Z=X_{L}\mathcal{W}_{L}$
6:return $Z$
$\bullet$ Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for GNN models with a large number of layers, information loss will still persist. Instead of adding the neighboring nodes, we will add a node, namely Cluster Nodes, representing the neighboring cluster, denoted by $\mathcal{C}_{G_{i}}$.
$\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$
Liu et al. (2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs. They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method after adding these Cluster Nodes and cross-cluster edges. In our work, we add cross-cluster edges; however, we do not sample from the constructed set of subgraphs.
Figure 2 illustrates appending these additional nodes using these approaches. The dashed line shows which nodes are added as part of these additional nodes. Predictions on these newly appended nodes will not contribute to loss calculation and weight update.
Lemma 1.
Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
Proof.
There are $2$ sets of nodes in $G_{i}$ as follows:
  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
Let $\mathcal{I}^{1}_{i}$ be the number of nodes whose information does not get passed on after $1$ layer of GNN for $G_{i}$.
$\mathcal{I}^{1}_{i}=\left|\bigcup_{v\in S_{2}}\mathcal{N}_{1}(v)-V(G_{i})\right|$
Now, |$\begin{split}\textup{Now, }\left|\bigcup_{v\in S_{1}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\right|=0\\ \textup{Also, }\mathcal{E}_{G_{i}}=\bigcup_{v\in S_{2}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\\ \end{split}$
$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
Hence when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
Algorithm 2 Training GNN (Train on $\mathcal{G}_{s}$)
1:$G=(V,E,X)$; Labels $Y$; Model $M$; Loss $\ell$;
2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$.
3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using P;
4:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$. Masks denote the nodes that are used for training;
5:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
6:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
7:$O\leftarrow[]$; $Y\leftarrow[]$;
8:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
9:     $O_{i}=M(A_{i},D_{i},X_{i})$;
10:     $O.append(O_{i}[mask_{i}])$;
11:     $Y.append(Y_{i}[mask_{i}])$;
12:end for
13:$Loss=\ell(O,Y)$;
14:Train $M$ to minimize $Loss$;
Using Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
$\bullet$ The number of nodes added to a subgraph via Cluster Nodes is less than or equal to the number of nodes added via the Extra Nodes approach. This is because, using Extra Nodes, we add all the neighboring nodes. However, we only add one representative node for each neighboring cluster using Cluster Nodes. As the GNN propagation depends on the number of nodes, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
$\bullet$ As mentioned before, the Extra Nodes approach reduces the information loss due to partitioning. However, with multiple layers of GNN, longer dependencies are not captured. In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Nodes’s representation as a combination of features of all the nodes present in the corresponding cluster. This results in sharing further node information in $1$-hop. Additionally, the transfer of information from one cluster to another is captured.
4.1 Node-level Task
We use Algorithm 1 with $L$ number of layers to create a standard GNN model for node-level tasks. Training the model on $G^{\prime}$ is similar to the Huang et al. (2021) training algorithm. We design Algorithm 2 to train the model on $\mathcal{G}_{s}$.
For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function. For the node regression task, we do not create any coarsened graph. Mean Absolute Error (MAE) is used as the loss function.
4.2 Graph-level Task
Unlike node-level tasks, the number of dataset samples is not decreasing. For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$. Each subgraph contains overlapping nodes when we use Extra Nodes or Cluster Nodes. Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Nodes or Cluster Nodes); otherwise, $False$.
Unlike node-level tasks, we cannot train and infer on $G^{\prime}$ and $\mathcal{G}_{s}$ using the same model. We use Algorithm 3 to create a model to train on $G^{\prime}$. We design Algorithm 4 to train the model on $\mathcal{G}_{s}$. We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task. We use Algorithm 2 without the backpropagation for inference.
Algorithm 3 Graph-Model-$G^{\prime}(L,A^{\prime},D^{\prime},X^{\prime})$
1:Number of Layers $L;A^{\prime};D^{\prime};X_{0}=X^{\prime}$
2:for $i=1$ to $L$ do
[…]
4:end for
5:$\bar{X}=\textup{MaxPooling}(X_{L})$
6:$Z=\bar{X}\mathcal{W}_{L}$
7:return $Z$
Algorithm 4 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L,mask)$
1:$\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$; Number of Layers $L$; $mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
2:$X\leftarrow[]$
3:for each subgraph $G_{i}\in\mathcal{G}_{s}$ do
4:     for $j=1$ to $L$ do
5:         $X_{j}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{j-1}\mathcal{W}_{j-1})$
6:     end for
7:     $X.append(X_{L}[mask_{i}])$
8:end for
9:$\bar{X}=\textup{MaxPooling}(X)$
10:$Z=\bar{X}\mathcal{W}_{L}$
11:return $Z$
Train Test
Classical $nd^{2}+n^{2}d$ $nd^{2}+n^{2}d$
Coarsening $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\begin{matrix}\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$
Table 1: This table shows the training and inference time complexity of FIT-GNN as compared to classical and coarsening approaches.
Train Test
Classical $n^{2}+nd+d^{2}$ $n^{2}+nd+d^{2}$
Coarsening $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $\begin{matrix}d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$
Table 2: This table shows the training and inference space complexity of FIT-GNN compared to classical and coarsening approaches.
4.3 Time and Space Complexity
The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X_{l}$ and $(d\times d)$ for $\mathcal{W}_{l}$. Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$. If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$. The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$. When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$.
Huang et al. (2021) improved the time and space complexity (mentioned in Table 1 and Table 2) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
Let us compare three different models. One is the classical model, where no coarsening or partitioning is done; the second is the Huang et al. (2021) approach; third is our approach where from the graph $G$ we create $G^{\prime}$ and $\mathcal{G}_{s}$.
The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph. Table 1 and Table 2 show the comparison between different approaches.
Lemma 2.
The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}$
Proof.
$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\\ =\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]\\ =k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ =nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ (\textup{where }k=nr)\\ =nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$.
Using this, we can say,
$\begin{split}&nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\\ \leq&nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d\\ =&nd^{2}+n^{2}d\end{split}$
Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$. ∎
For a given graph, if the above condition is satisfied, our approach has better inference time and space complexity than other approaches.
5 Experiments
Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
$\bullet$ Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$, then use the learned weight as an initialization for training and final inference on $\mathcal{G}_{s}$.
$\bullet$ Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$, then infer on $\mathcal{G}_{s}$.
$\bullet$ Gs-train-to-Gs-infer: Train and infer only on $\mathcal{G}_{s}$.
$\bullet$ Gc-train-to-Gc-infer: Unlike node-level tasks, we can use $G^{\prime}$ to infer for graph-level tasks. Therefore, in this setup, we train and infer on $G^{\prime}$.
For tasks like node regression, we can only perform Gs-train-to-Gs-infer because a coarsened graph is not created.
5.1 Experimental Setup
Dataset descriptions are in Section A of the appendix, and Section C of the appendix covers the hyperparameters, device configuration, and key packages used for the experiments. Source Code present at https://github.com/Roy-Shubhajit/FIT-GNN.
5.1.1 Node regression
We use three real-world network datasets for the node regression task: Chameleon, Crocodile, and Squirrel Rozemberczki et al. (2019). The nodes are divided into train (30%), validation (20%), and test (50%) sets. We perform 20 runs and report the average and standard deviation of the smallest 10 normalized MAE, with normalization achieved by dividing the loss by the standard deviation of the target values.
5.1.2 Node classification
We conducted node classification experiments on six real-world network datasets: Cora McCallum et al. (2000), Citeseer Giles et al. (1998), Pubmed Sen et al. (2008), Coauthor Physics Shchur et al. (2018), DBLP Tang et al. (2008), and a subset of OGBN-Products Hu et al. (2020) extracted using Leiden algorithm Traag et al. (2019). We used three types of splits for the datasets: we applied the public split from Yang et al. (2016) for Cora, Citeseer, and PubMed. In the “few” split, we set aside labeled nodes per class for training and validation, while the rest form the test set. We ran 20 trials for each configuration and reported the average and standard deviation of the top 10 accuracies.
r Dataset
Chameleon Crocodile Squirrel
Classical - $0.844\pm 0.000$ $0.853\pm 0.000$ $0.809\pm 0.000$
FIT-GNN $0.1$ $\mathbf{0.460\pm 0.002}$ $\mathbf{0.362\pm 0.002}$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.499\pm 0.006$ $0.364\pm 0.001$ $0.672\pm 0.003$
Table 3: This table shows the normalized MAE loss (lower is better) for node regression tasks on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
Fixed Few
r Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
$0.1$ $0.772\pm 0.006$ $\mathbf{0.829\pm 0.005}$ $0.676\pm 0.051$ $\mathbf{0.695\pm 0.016}$
$0.3$ $\mathbf{0.817\pm 0.016}$ $0.815\pm 0.005$ $0.694\pm 0.045$ $\mathbf{0.719\pm 0.011}$
$0.5$ $\mathbf{0.827\pm 0.001}$ $0.807\pm 0.005$ $0.688\pm 0.045$ $\mathbf{0.729\pm 0.007}$
$0.7$ $\mathbf{0.824\pm 0.002}$ $0.800\pm 0.008$ $0.679\pm 0.046$ $\mathbf{0.706\pm 0.012}$
Table 4: Results on Cora Dataset. The metrics shown here include the average and standard deviation of the top 10 accuracies over 20 runs. We show Gc-train-to-Gs-train experimental setup result for fixed and Gc-train-to-Gs-infer experimental result for few. The highest accuracy for each coarsening ratio is shown in bold.
$\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Dataset Split Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
Citeseer Few $0.583\pm 0.063$ $\mathbf{0.614\pm 0.021}$ $0.581\pm 0.052$ $\mathbf{0.624\pm 0.022}$
Fixed $\mathbf{0.711\pm 0.004}$ $0.677\pm 0.023$ $\mathbf{0.714\pm 0.003}$ $0.703\pm 0.007$
Pubmed Few $\mathbf{0.685\pm 0.052}$ $0.627\pm 0.116$ $\mathbf{0.687\pm 0.042}$ $0.676\pm 0.038$
Fixed $\mathbf{0.783\pm 0.005}$ $0.704\pm 0.022$ $\mathbf{0.784\pm 0.004}$ $0.677\pm 0.013$
DBLP Few $0.679\pm 0.056$ $\mathbf{0.683\pm 0.072}$ $0.648\pm 0.052$ $\mathbf{0.670\pm 0.034}$
Random $0.760\pm 0.021$ $\mathbf{0.789\pm 0.005}$ $0.745\pm 0.019$ $\mathbf{0.771\pm 0.007}$
Physics Coauthor Few $\mathbf{0.878\pm 0.036}$ $0.783\pm 0.028$ $\mathbf{0.908\pm 0.023}$ $0.877\pm 0.019$
Random $\mathbf{0.915\pm 0.014}$ $0.830\pm 0.048$ $\mathbf{0.934\pm 0.006}$ $0.910\pm 0.011$
OGBN-Products Random $0.296\pm 0.034$ $\mathbf{0.346\pm 0.014}$ $0.316\pm 0.016$ $\mathbf{0.406\pm 0.009}$
Table 5: Results for node classification tasks with accuracy as the metric (higher is better). We use the Cluster Nodes method to append additional nodes to subgraphs and Gc-train-to-Gs-train as experimental setup.
5.1.3 Graph Classification
Graph Classification has been conducted on the PROTEINS and AIDS dataset Morris et al. (2020); we randomly split the dataset into a $2:1:1$ ratio of train, validation, and test.
5.1.4 Graph Regression
We perform graph regression on the QM9 Wu et al. (2017), and a subset of ZINC Gómez-Bombarelli et al. (2016) datasets, using the same splits as in graph classification. Gilmer et al. (2017) grouped the $19$ properties of each molecule in QM9 dataset into four broad categories. We predict one property from each of the four broad categories. The four specific properties are the following: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
5.2 Result and Analysis
FIT-GNN
Dataset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (Subset) $0.688$ $\mathbf{0.578}$ $0.657$
QM9 ($\mu$) $0.869$ $\mathbf{0.841}$ $0.948$
QM9 ($\Delta\epsilon$) $1.012$ $\mathbf{0.875}$ $0.969$
QM9 ($\mathrm{ZPVE}$) $1.091$ $\mathbf{0.818}$ $0.840$
QM9 ($U^{\mathrm{ATOM}}$) $1.078$ $0.754$ $\mathbf{0.630}$
Table 6: Graph regression tasks results on ZINC (Subset) and QM9. The metrics shown are in terms of normalized MAE (lower is better) with Gs-train-to-Gs-infer experimental setup. The Cluster Nodes method is used to append additional nodes to subgraphs.
Gc-train-to
-Gs-infer
Gs-train-to
-Gs-infer
Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
AIDS $0.792$ $0.810$ $0.836$ $0.812$ $\mathbf{0.843}$
PROTEINS $0.623$ $0.636$ $0.681$ $0.812$ $\mathbf{0.821}$
Table 7: Results for the graph classification task on AIDS and PROTEINS are presented. The metrics reflect accuracy; higher values indicate better performance. The table also compares two experimental setups.
Cora Chameleon
Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
variation_neighborhoods $0.821\pm 0.003$ $0.803\pm 0.005$ $\mathbf{0.465\pm 0.005}$ $\mathbf{0.481\pm 0.006}$
algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
Table 8: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
Table 3 shows the node regression error and standard deviation on coarsening ratios $0.1$ and $0.3$. All the experiments for node regression are shown with Cluster node method. With a higher coarsening ratio, more subgraphs are created. As the coarsening ratio increases, the size of each subgraph also reduces, and important neighboring node information gets lost. Hence, the performance decreases with an increase in the coarsening ratio.
Table 4 shows the results for node classification accuracy and standard deviation on different coarsening ratios on the Cora dataset with Cluster Nodes approach. With coarsening ratios $0.1$ and $0.3$, the observed performance is better than other coarsening ratios, and the inference time taken is also less. This motivates us to experiment on other datasets with lower coarsening ratios. Table 5 shows the results for node classification on different datasets using coarsening ratios $0.1$ and $0.3$. From the results, it can be observed that the accuracy is comparable to the baseline. Since experimenting on the full OGBN-Products dataset is not feasible due to memory constraints, we create a proxy graph by taking the maximum-sized communities with the number of nodes that sum up to $165000$. For the subset of the OGBN-Products graph, the performance is better with the baseline even with $0.56\%$ training nodes. For larger datasets like Pubmed and Physics Coauthor, a lower coarsening ratio doesn’t perform better than the baseline.
Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset. All the results show that the FIT-GNN model’s performance is comparable to the baselines. It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high. It implies that finer subgraphs lose out global information about the molecule, which is necessary for the prediction.
Table 7 displays graph classification results using two experimental setups across two datasets. Training on coarsened graphs outperforms the baseline but discards important information. Conversely, training on subgraphs significantly improves predictions compared to the baseline.
PROTEINS ZINC (subset)
Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.629}$ $0.859$
algebraic_JC $\mathbf{0.783}$ $0.478$ $0.823$ $0.810$
kron $0.522$ $0.652$ $0.663$ $1.201$
heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
variation_edges $0.652$ $0.565$ $1.548$ $1.883$
variation_cliques $0.696$ $0.652$ $0.689$ $\mathbf{0.742}$
Table 9: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets. The metric for the PROTEINS dataset is accuracy (higher the better), and for the ZINC (subset) is normalized MAE (lower the better)
Table 8 and Table 9 show the comparison of performance with different coarsening algorithms used by Huang et al. (2021). We find that variation_neighborhoods is more consistent among all other coarsening algorithms, therefore we use it to report other results also.
5.3 Inference Time and Memory Comparison
FIT-GNN
Dataset Classical $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
Chameleon $0.0027$ $0.0016$ $\mathbf{0.0014}$
Squirrel $0.0081$ $0.0017$ $\mathbf{0.0014}$
Crocodile $0.0070$ $0.0015$ $\mathbf{0.0015}$
Cora $0.0026$ $\mathbf{0.0019}$ $0.0020$
Citeseer $0.0031$ $\mathbf{0.0018}$ $0.0019$
Pubmed $0.0042$ $0.0019$ $\mathbf{0.0018}$
DBLP $0.0063$ $0.0020$ $\mathbf{0.0018}$
Physics Coauthor $0.0252$ $0.0020$ $\mathbf{0.0017}$
OGBN-Products $0.1762$ $0.0017$ $\mathbf{0.0016}$
Table 10: Inference time (sec) comparison for a single node using standard GNN as a baseline and FIT-GNN with two different coarsening ratios. Lower is better. Here, we used Cluster Nodes.
Figure 3: Comparison of inference time memory consumption (in MB) for Coauthor Physics Dataset between baseline and FIT-GNN
FIT-GNN
Dastaset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
ZINC (subset) $\mathbf{0.00184}$ $\mathbf{0.00184}$ $0.00190$
QM9 $\mathbf{0.00173}$ $0.00180$ $0.00191$
AIDS $0.00163$ $\mathbf{0.00155}$ $0.00163$
PROTEINS $0.00165$ $\mathbf{0.00160}$ $0.00163$
Table 11: Inference time (sec) comparison for the graph-level task. We use the Cluster Nodes method to add additional nodes to each subgraph and Gc-train-to-Gc-infer experiment setup
As mentioned in Section 4.3, our inference time is lesser than the standard GNN if certain conditions are met. Therefore, we empirically show reduction of time and space by our FIT-GNN model during inference. For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
In Table 10, we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks. The baseline model processes the entire graph, increasing inference time, especially for large graphs. In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions. For larger datasets like the OGBN-Products, we see up to $\mathbf{100\times}$ reduction in inference time. Table 17 in the Appendix shows the detailed comparison of memory consumptions for different datasets for both Extra Nodes and Cluster Nodes method along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{70\times}$ less memory than the baseline.
Table 11 compares the inference time of graph classification and graph regression task. We randomly select $1000$ graphs from the test split and infer them for baseline and FIT-GNN. We also append additional nodes to each subgraph using Cluster Nodes method. The table shows how our method is comparable and sometimes faster than the baseline. Also, we observe that the inference time increases with a higher coarsening ratio. Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges. In table 16 in the Appendix, we show the estimated $\phi_{\textup{max}}$ value from the feature dimension of the nodes and the coarsening ratio. Combinations with negative numbers indicate the FIT-GNN-based approach won’t reduce the inference time.
Overall, the inference time and memory for all the tasks mentioned are drastically less than the classical approach while maintaining the performance mentioned in the previous section.
6 Conclusion
In this paper, we have focused on inference time and memory and presented a new way to utilize existing graph coarsening algorithms for GNNs. We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity. Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of the inference time and consuming a fraction of the memory. A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes.
Appendix A Dataset Description
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features Classes
PROTEINS $1113$ $19$ $72$ $3$ $2$
AIDS $2000$ $7$ $16$ $38$ $2$
Table 12: Summary of datasets used for Graph classification Dataset
Dataset
Number of
Graphs
Average
Nodes
Average
Edges
Features
Number of
Targets
QM9 $130831$ $8$ $18$ $11$ $19$
ZINC (subset) $10000$ $11$ $25$ $1$ $1$
Table 13: Summary of datasets used for Graph regression Dataset
Dataset Nodes Edges Features Classes
Cora $2708$ $5278$ $1433$ $7$
Citeseer $3327$ $4552$ $3703$ $6$
Pubmed $19717$ $44324$ $500$ $3$
DBLP $17716$ $52867$ $1639$ $4$
Physics
Coauthor
$34493$ $247962$ $8415$ $5$
OGBN-Products $165000$ $4340428$ $100$ $47$
Table 14: Summary of datasets used for Node classification Dataset
Dataset Nodes Edges Features
Number of
Targets
Chameleon $2277$ $31396$ $128$ $1$
Squirrel $5201$ $198423$ $128$ $1$
Crocodile $11631$ $170845$ $128$ $1$
Table 15: Summary of datasets used for Node regression Dataset
Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
PROTEINS $6.999$ $0.333$ $-1.000$ $-1.571$
AIDS $322.000$ $82.000$ $34.000$ $13.429$
QM9 $79.000$ $19.000$ $7.000$ $1.857$
ZINC (subset) $-11.000$ $-4.333$ $-3.000$ $-2.429$
Table 16: Calculated $\phi_{\textup{max}}$ values for different datasets and coarsening ratios as per Lemma 1.
Appendix B More on Extra Nodes and Cluster Nodes
B.1 Extra Nodes
There are $3$ sets of nodes in $G_{i}$.
  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})\right|$
When we use Extra Nodes, the information loss can be written as follows:
$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-\mathcal{E}_{G_{i}}\right|$
The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other. An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Node method is used.
B.2 Cluster Nodes
Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$. Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$. This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
Previously, according to Lemma 1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method. It was also easy to quantify in terms of the number of nodes. However, it is different for Cluster Nodes. Let us discuss the issues first.
$\bullet$ Only a weighted version of node information is shared with the subgraph. Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$. Then the information contributed by these nodes is $\frac{d(v_{a})x_{a}+d(v_{b})x_{b}}{\sum_{p}d(v_{p})}$. Here $d(v_{p})$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
$\bullet$ Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d(v_{p})x_{p}}{\sum_{p}d(v_{p})}$. This will capture further dependencies.
The performance of Cluster Node will depend on some distance or similarity metric between $x_{c},x_{d}$ and $f(C_{j})$.
Appendix C Experiment Details (Parameters and Device Configuration)
For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight. We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks. For both coarsened graph and set of subgraph based model, we set epochs to 300 and the number of layers of GCN to 2 for training. We set the hidden dimensions for each layer of GCN to $512$.
The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ $2.20$GHz, $256$GB RAM, NVIDIA A100 $40$GB GPU. We use Pytorch Geometric to to train our models.
Appendix D Results
This section shows the memory consumption for different datasets with different methods. Also, we show how the FIT-GNN model performs with different methods of appending additional nodes to subgraphs at different coarsening ratios.
FIT-GNN
Dataset Split $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$ Baseline
Cora cluster $1.310$ $0.648$ $0.694$ $0.868$ $15.743$
extra $2.072$ $1.134$ $1.017$ $1.017$
Citeseer cluster $31.559$ $2.148$ $1.253$ $1.686$ $49.488$
extra $31.559$ $3.114$ $1.788$ $1.788$
Pubmed cluster $0.613$ $0.569$ $0.571$ $0.573$ $41.227$
extra $0.798$ $0.590$ $0.590$ $0.590$
DBLP cluster $6.067$ $2.940$ $1.847$ $1.966$ $118.121$
extra $13.907$ $4.663$ $2.649$ $2.274$
Physics Coauthor cluster $24.025$ $13.088$ $13.388$ $14.353$ $1169.521$
extra $63.146$ $22.434$ $22.434$ $22.434$
Chameleon cluster $0.212$ $0.246$ $0.291$ $0.438$ $2.198$
extra $0.653$ $0.653$ $0.653$ $0.592$
Crocodile cluster $1.183$ $1.188$ $1.735$ $1.866$ $11.562$
extra $4.368$ $4.368$ $4.368$ $3.134$
Squirrel cluster $1.507$ $1.607$ $1.903$ $3.009$ $9.075$
extra $7.275$ $7.275$ $7.275$ $6.953$
OGBN-Products cluster $6.047$ $5.613$ $8.636$ $208.029$
Table 17: Summary of memory consumption of datasets used for node-level tasks. All units are in MegaBytes (MB)
r Extra Nodes Cluster Nodes
$0.1$ $0.798\pm 0.008$ $\mathbf{0.824\pm 0.003}$
$0.3$ $0.798\pm 0.008$ $\mathbf{0.807\pm 0.004}$
$0.5$ $0.807\pm 0.004$ $\mathbf{0.814\pm 0.004}$
$0.7$ $0.797\pm 0.007$ $\mathbf{0.808\pm 0.006}$
Table 18: Average of top 10 accuracies (higher the better) of node classification task on Cora dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the superior performance of the Cluster Nodes method over Extra Nodes.
r Extra Nodes Cluster Nodes
0.1 $0.504\pm 0.004$ $\mathbf{0.460\pm 0.003}$
0.3 $0.579\pm 0.006$ $\mathbf{0.499\pm 0.006}$
0.5 $0.585\pm 0.004$ $\mathbf{0.505\pm 0.004}$
0.7 $\mathbf{0.504\pm 0.006}$ $0.560\pm 0.004$
Table 19: Average of best 10 losses (lesser the better) of node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup. The table also demonstrates the better performance of the Cluster Nodes method over Extra Nodes.
Dataset Coarsening Ratio Top 10 Loss
Chameleon Baseline $0.844\pm 0.000$
$0.1$ $\mathbf{0.460\pm 0.002}$
$0.3$ $0.499\pm 0.006$
$0.5$ $0.587\pm 0.000$
$0.7$ $0.591\pm 0.080$
Crocodile Baseline $0.853\pm 0.000$
$0.1$ $\mathbf{0.362\pm 0.002}$
$0.3$ $0.364\pm 0.001$
$0.5$ $0.375\pm 0.001$
$0.7$ $0.384\pm 0.002$
Squirrel Baseline $0.809\pm 0.000$
$0.1$ $\mathbf{0.648\pm 0.002}$
$0.3$ $0.672\pm 0.003$
$0.5$ $0.729\pm 0.004$
$0.7$ $0.694\pm 0.004$
Table 20: Node Regression Results with Gs-train-to-Gs-infer experimental setup.

### v3

@@
-FIT-GNN: Faster Inference Time for GNNs Using Coarsening
+FIT-GNN: Faster Inference Time for GNNs that ‘FIT’ in Memory Using Coarsening
 Abstract
-Scalability of Graph Neural Networks (GNNs) remains a significant challenge, particularly when dealing with large-scale graphs.
-To tackle this, coarsening-based methods are used to reduce the graph into a smaller graph, resulting in faster computation.
+Scalability of Graph Neural Networks (GNNs) remains a significant challenge.
+To tackle this, methods like coarsening, condensation, and computation trees are used to train on a smaller graph, resulting in faster computation.
 Nonetheless, prior research has not adequately addressed the computational costs during the inference phase.
-This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during both training and inference phases.
-We demonstrate two different methods (Extra-Nodes and Cluster-Nodes).
-Our study also proposes a unique application of the coarsening algorithm for graph-level tasks, including graph classification and graph regression, which have not yet been explored.
-We conduct extensive experiments on multiple benchmark datasets in the order of $100K$ nodes to evaluate the performance of our approach.
-The results demonstrate that our method achieves competitive performance in tasks involving classification and regression on nodes and graphs, compared to traditional GNNs, while having single-node inference times that are orders of magnitude faster.
-Furthermore, our approach significantly reduces memory consumption, allowing training and inference on low-resource devices where traditional methods struggle.
-$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally to this work
-[…]
-images/img1
-Figure 1: The figure shows how the coarsening algorithm reduces a graph $G$ with $15$ nodes to a coarsened graph $G^{\prime}$ with $3$ nodes. It also shows which nodes belong to which partition.
+This paper presents a novel approach to improve the scalability of GNNs by reducing computational burden during the inference phase using graph coarsening.
+We demonstrate two different methods – Extra Nodes and Cluster Nodes.
+Our study extends the application of graph coarsening for graph-level tasks, including graph classification and graph regression.
+We conduct extensive experiments on multiple benchmark datasets to evaluate the performance of our approach.
+Our results show that the proposed method achieves orders of magnitude improvements in single-node inference time compared to traditional approaches.
+Furthermore, it significantly reduces memory consumption for node and graph classification and regression tasks, enabling efficient training and inference on low-resource devices where conventional methods are impractical.
+Notably, these computational advantages are achieved while maintaining competitive performance relative to baseline models.
 1 Introduction
-Graph Neural Networks (GNNs) have proven to be a tool of surprising versatility and modeling capacity.
-However, there are still a few issues that limit their general applicability, the chief among these being the scalability of the GNN models.
-This scalability issue is problematic for both the training and the inference phases.
-Previous research has attempted to address this by reducing the graph into a smaller graph in order to reduce the computational cost.
-One such broad technique is graph coarsening methods, which judiciously remove nodes and edges from the graph.
-However, there are still certain points of weakness of these algorithms.
-For one, coarsening methods can potentially remove important node information.
-Secondly, and more pertinent to the current discussion, coarsening methods only reduce the training computation; they do not address the inference time.
-The next obvious question would be to ask why we cannot use coarsening methods for inference as well.
-Given a graph $G$, coarsening algorithms create a coarsened graph $G^{\prime}$.
-While $G^{\prime}$ is smaller than $G$, the nodes in $G^{\prime}$ are combinations of nodes in $G$, which can be both train and test nodes.
-As a result, the prediction per test node is not available.
-In another scenario, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes.
-Also, let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively.
-The coarsening algorithm will take the majority label and assign class 00 to $v^{\prime}$.
-Hence, the model will be trained to predict only the class 00, leading to discarding the quantification of the model’s performance on predicting less represented nodes.
-This is the research gap that we seek to address in this work.
-We use coarsening algorithms to partition the graph into a set of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$.
-We show methods to train on these subgraphs that show equivalent results to non-coarsened baselines and sometimes perform better.
-This comes hand in hand with a significant reduction in inference time and memory consumption.
-We demonstrate two different methods to append additional nodes in the subgraphs— Extra Nodes and Cluster Nodes that respectively append additional nodes to each subgraph, either by adding neighboring nodes, or by adding representational nodes corresponding a neighboring cluster.
-We theoretically characterize the upper bound of the coarsening ratio for which the time complexity of our approach is better than the naive uncoarsened approach.
-In node classification and node regression tasks, the baselines typically require the entire dataset during inference to get a prediction for a single node.
-In contrast, our method only needs to pass the subgraph to which the node belongs.
-Our results show up to $\mathbf{100\times}$ reduction in inference time and up to $\mathbf{70\times}$ less memory than the baseline while achieving equivalent or better performance.
+Graph Neural Networks (GNNs) have demonstrated remarkable versatility and modeling capabilities.
+However, several challenges still hinder their widespread applicability, with scalability being the most significant concern.
+This scalability issue affects both the training and inference phases.
+Previous research has sought to mitigate the higher training computation by using three broad categories of methods: 1.
+Training every iteration on the sampled subgraph 2.
+Training on a coarsened graph generated from a coarsening algorithm 3.
+Training on a synthetic graph that mimics the original graph .
+Unfortunately, these smaller graphs can only be utilized for training purposes and have not alleviated the computational burden during inference since the method mentioned above requires the whole graph during the inference phase.
+Consequently, for larger graphs, not only do inference times increase, but memory consumption may also exceed limitations.
+To mitigate the high computational cost associated with the inference phase, we decompose the original graph $G$ into a collection of subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$.
+This transformation yields substantial improvements in inference efficiency, achieving up to a $\mathbf{100\times}$ speedup for single-node prediction.
+Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges.
+We provide a theoretical characterization of the upper bound on the coarsening ratio under which our method’s time and space complexity remain superior to that of the standard approach.
+Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
+While partitioning the graphs into subgraphs seems to be a naive transformation, it comes with the cost of information loss corresponding to the edges removed.
+We overcome this loss by appending additional nodes in the subgraphs using two different methods – Extra Nodes and Cluster Nodes that respectively append neighboring nodes or representational nodes corresponding to neighboring clusters.
+These additional nodes resulted in equivalent or better performance in terms of accuracy and error compared to baselines.
+Our contributions are summarized as follows:
+  1. 1.
+    We reduce inference time and memory by applying graph coarsening to partition graphs into subgraphs, mitigating information loss through Extra Node and Cluster Node strategies.
+  2. 2.
+    We provide a theoretical analysis of our approach, establishing a lower inference computational complexity compared to baselines.
+  3. 3.
+    Our method scales well to extremely large graphs that baselines fail to process, while consistently delivering strong performance on 13 real-world datasets.
+Figure 1: The figure shows the overall pipeline of our proposed method and its comparison with traditional training and inference of GNNs. This pipeline is made for node-level tasks.
+2 Related Work
+The aim of reducing computational costs has been approached through various methods.
+Research by Chiang et al.
+[1] and Zeng et al.
+[2] uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training.
+Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen and Zhu [3], Chen et al.
+[4], Cong et al.
+[5], Zou et al.
+[6].
+Additionally, Fahrbach et al.
+[7] introduced a coarsening algorithm based on Schur complements to create embeddings for vertices in large graphs, while Huang et al.
+[8] (SGGC) applied GNNs to coarsened graphs, reducing nodes and edges to lower training time.
+Kumar et al.
+[9] proposed feature-based coarsening algorithms.
+Jin et al.
+[10] (GCOND), Wang et al.
+[11] (GC-SNTK) presented algorithms to generate a synthetic graph that mimics the original graph.
+However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden.
+Gupta et al.
+[12] (BONSAI) addressed this issue by looking into computational trees and training on just the relevant computational trees.
+However, these approaches do not decrease inference time since the inference phase takes the whole graph as input.
+Later, Xue et al.
+[13] shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach.
+In the graph-level task, Jin et al.
+[14] (DOSCOND) and Xu et al.
+[15] (KIDD) tackled the graph classification problem following up on the approach of GCOND and GC-SNTK.
+However, the graph regression problem was not explored in this direction.
 3 Preliminaries
@@
 Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix.
-The feature of each node $v_{i}$ in the graph is the $i$th row vector in $X$ represented as $x_{i}\in\mathbb{R}^{d}$.
 Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges.
-Let us assume $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$.
+Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$.
 $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$.
@@
 Graph Neural Network is a neural network designed to work for graph-structured data.
-The representation of each node in the graph is updated recursively by aggregation and transforming node representations of its neighbors.
-Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
-$X_{l+1}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X_{l}\mathcal{W}_{l})$ (1)
-where $X_{l}$ is the node representation after $l$ layers of GCN, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is an identity matrix of same dimension as $A$.
-$\mathcal{W}_{l}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
+The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors.
+Kipf and Welling [16] introduced Graph Convolutional Network (GCN) as follows:
+$X^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(l)}\mathcal{W}^{(l)})$ (1)
+where $X^{(l)}$ is the node representation after $l$ layers, $\tilde{A}=A+I$, $\tilde{D}=D+I$, $I$ is the identity matrix of same dimension as $A$.
+$\mathcal{W}^{(l)}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
 3.3 Graph Coarsening and Graph Partitioning
-Loukas (2018) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
+Loukas [17] discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
 We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes.
@@
 Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$.
-For the coarsened graph, the adjacency matrix $A^{\prime}=P^{T}AP$.
+$A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$.
 Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$.
-Huang et al.
-(2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{j,j}=|C_{j}|,j=1,2,\dots,k$.
-Along with a coarsened graph, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$.
+SGGC used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
+Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$.
 The number of nodes in each partition $C_{i}$ is denoted by $n_{i}$.
@@
 $A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
-3 Related Work
-The goal of reducing computation costs has been tackled via different methodologies.
-Research like Chiang et al.
-(2019) and Zeng et al.
-(2020) use a subgraph sampling-based approach.
-They sample a small subgraph at each training iteration and train the GNN.
-Other techniques like layer-wise sampling along with mini-batch training have been studied by  Chen et al.
-(2018); Chen and Zhu (2018); Cong et al.
-(2020); Zou et al.
-(2019).
-Fahrbach et al.
-(2020) introduced a new coarsening algorithm based on Schur complements to generate embeddings for vertices in a large graph.
-Huang et al.
-(2021) introduced using GNN for a coarsened graph.
-The motivation was to utilize graph coarsening algorithms to reduce the number of nodes and edges in the graph, resulting in less training time.
-This approach helps to train GNN for graphs with a large number of nodes.
-However, inference time is not reduced because the nodes cannot be removed during inference.
-This training process also caused the removal of important node-level information.
-Later, Xue et al.
-(2022) explored shifting the training process from graph-level training to subgraph-level.
-Coarsening algorithms were used to split the graph into $k-$subgraphs, resulting in the removal of a few edges.
-Further, to speed up the training process, they propose a multi-device training approach where subgraphs are sent to different devices.
-This results in a faster training process.
-Kumar et al.
-(2023) further proposed the usage of feature-based coarsening algorithms.
+[…]
+images/img2
+Figure 2: Methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
 4 Methodology
-In Huang et al.
-(2021), a graph $G=(V,E,X,W)$ is reduced to $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ using a partition matrix $P$ generated from the coarsening algorithm.
+In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm.
 Here $X^{\prime}=P^{T}X$.
 The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task.
-While using $\arg\max$, some essential information is lost.
-Therefore, we follow a subgraph-level training process where we don’t remove these label information.
-To explore another node-level task, such as node regression, functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets.
-Hence, our goal is to use subgraphs for training and inference purposes.
-To tackle the issue of loss of information corresponding to the nodes (and their edges) present at the periphery of the subgraphs after partitioning the graph, we append nodes in two ways:
-[…]
-images/img2
-Figure 2: This figure shows the methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
-$\bullet$ Extra Nodes: Adding the $1$-hop neighbouring nodes in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
-$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$
-However, we don’t perform backpropagation of the prediction on $\mathcal{E}_{G_{i}}$.
+Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information.
+While this approach is novel, essential label information is lost.
+For instance, suppose a node $v^{\prime}\in G^{\prime}$ is a combination of $v_{i},v_{j},v_{k}\in G$ where $v_{i},v_{j},v_{k}$ are test nodes.
+Let us assume that the classes of $v_{i},v_{j},v_{k}$ are $0,0,2$ respectively.
+$\arg\max$ will take the majority label and assign class 0 to $v^{\prime}$.
+Hence, the model will be trained to predict only the class 0, leading to the discarding of the quantification of the model’s performance on predicting less represented nodes.
+Similarly, to explore another node-level task, such as node regression, aggregation functions such as the mean of regression targets wouldn’t be appropriate because they lose the variance of the targets.
+To mitigate label information loss, we use original label information without any aggregation function like $\arg\max$ or mean, thus motivating us to follow a subgraph-level training process.
+However, partitioning the graph into subgraphs leads to a loss of neighborhood information at the periphery of subgraphs. To tackle this, we append nodes in two ways after partitioning:
+  • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
 Xue et al.
-(2022) introduced this approach to reduce information loss after partitioning.
-An edge between two Extra Nodes in $\mathcal{E}_{G_{i}}$ is added if they were connected in $G$.
-Algorithm 1 Node-Model$(L,A,D,X)$
-1:Number of Layers $L$; $A$; $D$; $X_{0}=X$;
-2:for $i=1$ to $L$ do
-[…]
-4:end for
-5:$Z=X_{L}\mathcal{W}_{L}$
-6:return $Z$
-$\bullet$ Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for GNN models with a large number of layers, information loss will still persist.
-Instead of adding the neighboring nodes, we will add a node, namely Cluster Nodes, representing the neighboring cluster, denoted by $\mathcal{C}_{G_{i}}$.
-$\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$
+[13] introduced this approach to reduce information loss after partitioning.
+We add an edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
+  • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist.
+Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$.
 Liu et al.
-(2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs.
-They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method after adding these Cluster Nodes and cross-cluster edges.
-In our work, we add cross-cluster edges; however, we do not sample from the constructed set of subgraphs.
+[18] introduced this method to add nodes in the graph instead of adding to separate subgraphs.
+They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph.
+In our work, we add cross-cluster edges but do not sample from the constructed set of subgraphs.
+$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\quad\text{and}\quad\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (2)
 Figure 2 illustrates appending these additional nodes using these approaches.
 The dashed line shows which nodes are added as part of these additional nodes.
-Predictions on these newly appended nodes will not contribute to loss calculation and weight update.
-Lemma 1.
+For subgraph-level training, specifically for node-level tasks, the newly appended nodes do not contribute to the weight update as the loss is not backpropagated based on the predictions of these nodes.
+This is done since the appended nodes might contain information from the test nodes.
+Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Node or Cluster Node) and are training nodes; otherwise, $False$.
+Lemma 4.1.
 Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
 Proof.
-There are $2$ sets of nodes in $G_{i}$ as follows:
-  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
-  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
-Let $\mathcal{I}^{1}_{i}$ be the number of nodes whose information does not get passed on after $1$ layer of GNN for $G_{i}$.
-$\mathcal{I}^{1}_{i}=\left|\bigcup_{v\in S_{2}}\mathcal{N}_{1}(v)-V(G_{i})\right|$
-Now, |$\begin{split}\textup{Now, }\left|\bigcup_{v\in S_{1}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\right|=0\\ \textup{Also, }\mathcal{E}_{G_{i}}=\bigcup_{v\in S_{2}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\\ \end{split}$
-$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
-Hence when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
+In Appendix A.1. ∎
+Using the Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
+  • From Equation 2, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$ This is because, using Extra Nodes, we add all the neighboring nodes.
+However, using Cluster Nodes, we only add one representative node for each neighboring cluster.
+Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
+  • Lemma 4.1 shows that the Extra Nodes approach reduces the information loss due to partitioning for $1$ Layer of GNN.
+However, with multiple layers of GNN, longer dependencies are not captured.
+In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Node’s representation as a combination of features of all the nodes present in the corresponding cluster.
+This results in sharing further node information in $1$-hop.
+Additionally, the transfer of information from one cluster to another is captured.
 Algorithm 1 Training GNN (Train on $\mathcal{G}_{s}$)
-1:$G=(V,E,X)$; Labels $Y$; Model $M$; Loss $\ell$;
+1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
 2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$.
 3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using $P$;
-4:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$. Masks denote the nodes that are used for training;
-5:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
-6:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
-7:$O\leftarrow[]$; $Y\leftarrow[]$;
-8:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
-9:     $O_{i}=M(A_{i},D_{i},X_{i})$;
-10:     $O.append(O_{i}[mask_{i}])$;
-11:     $Y.append(Y_{i}[mask_{i}])$;
-12:end for
-13:$Loss=\ell(O,Y)$;
-14:Train $M$ to minimize $Loss$;
-Using Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
-$\bullet$ The number of nodes added to a subgraph via Cluster Nodes is less than or equal to the number of nodes added via the Extra Nodes approach.
-This is because, using Extra Nodes, we add all the neighboring nodes.
-However, we only add one representative node for each neighboring cluster using Cluster Nodes.
-As the GNN propagation depends on the number of nodes, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
-$\bullet$ As mentioned before, the Extra Nodes approach reduces the information loss due to partitioning.
-However, with multiple layers of GNN, longer dependencies are not captured.
-In contrast, the Cluster Nodes approach overcomes this by computing a single Cluster Nodes’s representation as a combination of features of all the nodes present in the corresponding cluster.
-This results in sharing further node information in $1$-hop.
-Additionally, the transfer of information from one cluster to another is captured.
+4:if Append additional nodes then
+[…]
+6:end if
+7:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
+8:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
+9:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
+[…]
+11:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
+12:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\quad\#$ Algorithm 4
+13:  $O.append(O_{i}[mask_{i}])$;
+14:  $Y.append(Y_{i}[mask_{i}])$;
+15:end for
+[…]
+17:Train $M$ to minimize LossLoss;
 4.1 Node-level Task
-We use Algorithm 1 with $L$ number of layers to create a standard GNN model for node-level tasks.
-Training the model on $G^{\prime}$ is similar to the Huang et al.
-(2021) training algorithm.
-We design Algorithm 2 to train the model on $\mathcal{G}_{s}$.
+We use Algorithm 4 with $L$ number of layers to create a standard GNN model $M$ for node-level tasks.
+While we use Algorithm 3 to train a model on the coarsened graph $G^{\prime}$, we propose Algorithm 1 to train a model on the set of subgraphs $\mathcal{G}_{s}$.
+Algorithm 3 utilizes the partition matrix $P$ generated by the coarsening algorithm to construct and train on the coarsened graph $G^{\prime}$ and the coarsened label information $Y^{\prime}$.
+On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$.
 For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function.
-For the node regression task, we do not create any coarsened graph.
+For the node regression task, we do not create $G^{\prime}$.
 Mean Absolute Error (MAE) is used as the loss function.
 4.2 Graph-level Task
-Unlike node-level tasks, the number of dataset samples is not decreasing.
 For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$.
-Each subgraph contains overlapping nodes when we use Extra Nodes or Cluster Nodes.
-Therefore, a boolean array $mask_{i}$ is created for each subgraph $G_{i}$ such that $True$ is set for nodes that actually belong to the subgraph $G_{i}$ (not as an Extra Nodes or Cluster Nodes); otherwise, $False$.
-Unlike node-level tasks, we cannot train and infer on $G^{\prime}$ and $\mathcal{G}_{s}$ using the same model.
-We use Algorithm 3 to create a model to train on $G^{\prime}$.
-We design Algorithm 4 to train the model on $\mathcal{G}_{s}$.
+Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model to train on $G^{\prime}$ and Algorithm 2 to create a model to train on $\mathcal{G}_{s}$.
 We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error(MAE) for the graph regression task.
-We use Algorithm 2 without the backpropagation for inference.
-Algorithm 3 Graph-Model-$G^{\prime}(L,A^{\prime},D^{\prime},X^{\prime})$
-1:Number of Layers $L;A^{\prime};D^{\prime};X_{0}=X^{\prime}$
-2:for $i=1$ to $L$ do
+Algorithm 2 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L)$
 […]
-4:end for
-5:$\bar{X}=\textup{MaxPooling}(X_{L})$
-6:$Z=\bar{X}\mathcal{W}_{L}$
-7:return $Z$
-Algorithm 4 Graph-Model-$\mathcal{G}_{s}(\mathcal{G}_{s},L,mask)$
-1:$\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$; Number of Layers $L$; $mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
-2:$X\leftarrow[]$
+[…]
 3:for each subgraph $G_{i}\in\mathcal{G}_{s}$ do
 4:  for $j=1$ to $L$ do
-5:         $X_{j}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{j-1}\mathcal{W}_{j-1})$
+[…]
 6:  end for
-7:     $X.append(X_{L}[mask_{i}])$
+[…]
 8:end for
-9:$\bar{X}=\textup{MaxPooling}(X)$
-10:$Z=\bar{X}\mathcal{W}_{L}$
+[…]
+[…]
 11:return $Z$
-Train Test
+4.3 Time and Space Complexity
+(a) Time Complexity
+Train Inference (Full Graph)
 Classical $nd^{2}+n^{2}d$ $nd^{2}+n^{2}d$
-Coarsening $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
-FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\begin{matrix}\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$
-Table 1: This table shows the training and inference time complexity of FIT-GNN as compared to classical and coarsening approaches.
-Train Test
+SGGC $kd^{2}+k^{2}d$ $nd^{2}+n^{2}d$
+FIT-GNN $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$
+(b) Space Complexity
+Train Inference
 Classical $n^{2}+nd+d^{2}$ $n^{2}+nd+d^{2}$
-Coarsening $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
-FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $\begin{matrix}d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$
-Table 2: This table shows the training and inference space complexity of FIT-GNN compared to classical and coarsening approaches.
-4.3 Time and Space Complexity
-The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X_{l}$ and $(d\times d)$ for $\mathcal{W}_{l}$.
+SGGC $k^{2}+kd+d^{2}$ $n^{2}+nd+d^{2}$
+FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $d^{2}+\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]$
+Table 1: Training and inference time and space complexity of FIT-GNN compared to classical and SGGC approaches.
+The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X^{(l)}$ and $(d\times d)$ for $\mathcal{W}^{(l)}$.
 Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$.
@@
 When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$.
-Huang et al.
-(2021) improved the time and space complexity (mentioned in Table 1 and Table 2) for training the network by reducing the number of nodes from $n$ to $k$.
-However, the inference time and space complexity remain the same.
+The complexity is based on GCN.
+For other architectures like GAT [19], the complexity changes.
+SGGC improved the time and space complexity (Table 1) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
 Let us compare three different models.
-One is the classical model, where no coarsening or partitioning is done; the second is the Huang et al.
-(2021) approach; third is our approach where from the graph $G$ we create $G^{\prime}$ and $\mathcal{G}_{s}$.
+One is the classical model, where no coarsening or partitioning is done; the second is SGGC, where training is performed on a smaller graph $G^{\prime}$ and inference on $G$; the third is our approach, FIT-GNN, where both training and inference are performed on $\mathcal{G}_{s}$.
 The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph.
-Table 1 and Table 2 show the comparison between different approaches.
-Lemma 2.
-The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}$
+Table 1 compares different approaches.
+Lemma 4.2.
+The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}.$
 Proof.
-$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\\ =\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]\\ =k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ =nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\\ (\textup{where }k=nr)\\ =nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
+$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]$
+$=k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d$
+$=nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\quad(\textup{where }k=nr)$
+$=nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
 Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$.
 Using this, we can say,
-$\begin{split}&nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\\ \leq&nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d\\ =&nd^{2}+n^{2}d\end{split}$
+$nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\leq nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d=nd^{2}+n^{2}d\\$
 Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$. ∎
-For a given graph, if the above condition is satisfied, our approach has better inference time and space complexity than other approaches.
+For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches.
+In Section C, we show that, for datasets with millions of nodes, the conditions outlined in Lemma 4.2 can indeed be satisfied.
+As the number of nodes increases, the gap in inference time between baseline methods and FIT-GNN widens.
+In table 9, we show the upper bound of $\phi_{\max}$ calculated as per the conditions mentioned in Lemma 4.2 for different datasets corresponding to different coarsening ratios.
+Negative upper bound is observed for some datasets, implying that our method won’t give better computational complexities in those scenarios.
 5 Experiments
+Let training on the coarsened graph $G^{\prime}$ be referred to as Gc-train, subgraph-level training as Gs-train, and subgraph-level inference process as Gs-infer.
 Once we have constructed $G^{\prime}$ and $\mathcal{G}_{s}$, we train FIT-GNN in 4 different setups:
-$\bullet$ Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$, then use the learned weight as an initialization for training and final inference on $\mathcal{G}_{s}$.
-$\bullet$ Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$, then infer on $\mathcal{G}_{s}$.
-$\bullet$ Gs-train-to-Gs-infer: Train and infer only on $\mathcal{G}_{s}$.
-$\bullet$ Gc-train-to-Gc-infer: Unlike node-level tasks, we can use $G^{\prime}$ to infer for graph-level tasks. Therefore, in this setup, we train and infer on $G^{\prime}$.
-For tasks like node regression, we can only perform Gs-train-to-Gs-infer because a coarsened graph is not created.
-5.1 Experimental Setup
-Dataset descriptions are in Section A of the appendix, and Section C of the appendix covers the hyperparameters, device configuration, and key packages used for the experiments.
-Source Code present at https://github.com/Roy-Shubhajit/FIT-GNN.
-5.1.1 Node regression
-We use three real-world network datasets for the node regression task: Chameleon, Crocodile, and Squirrel Rozemberczki et al.
-(2019).
-The nodes are divided into train (30%), validation (20%), and test (50%) sets.
-We perform 20 runs and report the average and standard deviation of the smallest 10 normalized MAE, with normalization achieved by dividing the loss by the standard deviation of the target values.
-5.1.2 Node classification
-We conducted node classification experiments on six real-world network datasets: Cora McCallum et al.
-(2000), Citeseer Giles et al.
-(1998), Pubmed Sen et al.
-(2008), Coauthor Physics Shchur et al.
-(2018), DBLP Tang et al.
-(2008), and a subset of OGBN-Products Hu et al.
-(2020) extracted using Leiden algorithm Traag et al.
-(2019).
-We used three types of splits for the datasets: we applied the public split from Yang et al.
-(2016) for Cora, Citeseer, and PubMed.
-In the “few” split, we set aside labeled nodes per class for training and validation, while the rest form the test set.
-We ran 20 trials for each configuration and reported the average and standard deviation of the top 10 accuracies.
-r Dataset
-Chameleon Crocodile Squirrel
-Classical - $0.844\pm 0.000$ $0.853\pm 0.000$ $0.809\pm 0.000$
-FIT-GNN $0.1$ $\mathbf{0.460\pm 0.002}$ $\mathbf{0.362\pm 0.002}$ $\mathbf{0.648\pm 0.002}$
-$0.3$ $0.499\pm 0.006$ $0.364\pm 0.001$ $0.672\pm 0.003$
-Table 3: This table shows the normalized MAE loss (lower is better) for node regression tasks on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
-Fixed Few
-r Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
-$0.1$ $0.772\pm 0.006$ $\mathbf{0.829\pm 0.005}$ $0.676\pm 0.051$ $\mathbf{0.695\pm 0.016}$
-$0.3$ $\mathbf{0.817\pm 0.016}$ $0.815\pm 0.005$ $0.694\pm 0.045$ $\mathbf{0.719\pm 0.011}$
-$0.5$ $\mathbf{0.827\pm 0.001}$ $0.807\pm 0.005$ $0.688\pm 0.045$ $\mathbf{0.729\pm 0.007}$
-$0.7$ $\mathbf{0.824\pm 0.002}$ $0.800\pm 0.008$ $0.679\pm 0.046$ $\mathbf{0.706\pm 0.012}$
-Table 4: Results on Cora Dataset.
-The metrics shown here include the average and standard deviation of the top 10 accuracies over 20 runs.
-We show Gc-train-to-Gs-train experimental setup result for fixed and Gc-train-to-Gs-infer experimental result for few.
-The highest accuracy for each coarsening ratio is shown in bold.
-$\mathbf{r=0.1}$ $\mathbf{r=0.3}$
-Dataset Split Huang et al. (2021) FIT-GNN Huang et al. (2021) FIT-GNN
-Citeseer Few $0.583\pm 0.063$ $\mathbf{0.614\pm 0.021}$ $0.581\pm 0.052$ $\mathbf{0.624\pm 0.022}$
-Fixed $\mathbf{0.711\pm 0.004}$ $0.677\pm 0.023$ $\mathbf{0.714\pm 0.003}$ $0.703\pm 0.007$
-Pubmed Few $\mathbf{0.685\pm 0.052}$ $0.627\pm 0.116$ $\mathbf{0.687\pm 0.042}$ $0.676\pm 0.038$
-Fixed $\mathbf{0.783\pm 0.005}$ $0.704\pm 0.022$ $\mathbf{0.784\pm 0.004}$ $0.677\pm 0.013$
-DBLP Few $0.679\pm 0.056$ $\mathbf{0.683\pm 0.072}$ $0.648\pm 0.052$ $\mathbf{0.670\pm 0.034}$
-Random $0.760\pm 0.021$ $\mathbf{0.789\pm 0.005}$ $0.745\pm 0.019$ $\mathbf{0.771\pm 0.007}$
-Physics Coauthor Few $\mathbf{0.878\pm 0.036}$ $0.783\pm 0.028$ $\mathbf{0.908\pm 0.023}$ $0.877\pm 0.019$
-Random $\mathbf{0.915\pm 0.014}$ $0.830\pm 0.048$ $\mathbf{0.934\pm 0.006}$ $0.910\pm 0.011$
-OGBN-Products Random $0.296\pm 0.034$ $\mathbf{0.346\pm 0.014}$ $0.316\pm 0.016$ $\mathbf{0.406\pm 0.009}$
+  • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$.
+  • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights.
+  • Gs-train-to-Gs-infer: Perform subgraph-level training and inference on $\mathcal{G}_{s}$.
+  • Gc-train-to-Gc-infer: Unlike node-level tasks, we can infer on $G^{\prime}$ (Gc-infer) itself for graph-level tasks after training on $G^{\prime}$.
+Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
+All four setups can be applied to graph-level tasks.
+For the node classification task, we apply all setups except Gc-train-to-Gc-infer.
+For the node regression task, we only perform Gs-train-to-Gs-infer because a coarsened graph is not created for node regression datasets.
+[…]
+Tasks Datasets
+Node Regression
+(Train=30%, Val=20%, Test=50%)
+Chameleon [20]
+Crocodile [20]
+Squirrel [20]
+Node Classification
+(Train=20*c, Val=30*c, Test=Rest)
+Cora [21]
+Citeseer [22]
+Pubmed [23]
+Coauthor Physics [24]
+DBLP [25]
+OGBN-Products [26]
+Graph Regression
+(Train=50%, Val=25%, Test=25%)
+QM9 [27]
+ZINC (Subset) [28]
+Graph Classification
+(Train=50%, Val=25%, Test=25%)
+PROTEINS [29]
+AIDS [29]
+Table 2: The table shows the datasets used and the splits in each dataset.
+‘c’ represents the number of classes for the dataset.
+We applied the public split “fixed” from [30] for Cora, Citeseer, and PubMed.
+For the “random” split in node classification, we keep 2020 nodes per class for training, $30$ nodes for validation, and the rest as test nodes.
+[31] grouped the 1919 properties of each molecule in QM9 dataset into four broad categories.
+We predict one property from each broad category: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
+6 Result and Analysis
+Methods Models
+Reduction
+Ratio (r)
+Dataset
+Chameleon Crocodile Squirel
+Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
+GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
+SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
+GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
+FIT-GNN GCN 0.1 $\mathbf{0.496\pm 0.005}$ $0.364\pm 0.001$ $0.663\pm 0.003$
+GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
+SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
+GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
+Table 3: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
+variation_neighborhoods coarsening algorithm is used.
+Methods Models
+Reduction
+Ratio (r)
+Dataset
+Cora Citeseer Pubmed DBLP Physics
+Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $\mathbf{0.933\pm 0.000}$
+GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
+SGGC GCN 0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
+0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
+GAT 0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
+0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
+GCOND GCN 0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
+0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
+GAT 0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
+0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
+BONSAI GCN 0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
+0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
+GAT 0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
+0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
+FIT-GNN GCN 0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $0.932\pm 0.000$
+0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
+GAT 0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
+0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
 Table 4: Results for node classification tasks with accuracy as the metric (higher is better).
-We use the Cluster Nodes method to append additional nodes to subgraphs and Gc-train-to-Gs-train as experimental setup.
-5.1.3 Graph Classification
-Graph Classification has been conducted on the PROTEINS and AIDS dataset Morris et al. (2020); we randomly split the dataset into a $2:1:1$ ratio of train, validation, and test.
-5.1.4 Graph Regression
-We perform graph regression on the QM9 Wu et al.
-(2017), and a subset of ZINC Gómez-Bombarelli et al.
-(2016) datasets, using the same splits as in graph classification.
-Gilmer et al.
-(2017) grouped the $19$ properties of each molecule in QM9 dataset into four broad categories.
-We predict one property from each of the four broad categories.
-The four specific properties are the following: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
-5.2 Result and Analysis
+We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
+variation_neighborhoods coarsening algorithm is used.
+6.1 Model Performance
+Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$
+Full $0.873\pm 0.000$
+SGGC OOM
+GCOND OOM
+BONSAI OOM
+FIT-GNN $\mathbf{0.894\pm 0.000}$
+Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE) with coarsening ratios $0.1$.
+Results using different coarsening ratios are in Table 12.
+As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost.
+Hence, an increase in the coarsening ratio decreases performance.
+Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach.
+Comparison with other coarsening ratios is in Table 13.
+Results show that the accuracy is comparable to the baselines.
+Table 5 shows results on the OGBN-Products dataset.
+Result on training on the full dataset from the OGBN-Products leaderboard [Link].
+For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing.
+For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
+Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset.
+More results at Table 14.
+All the results show that the FIT-GNN model’s performance is better than the baselines.
+It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high.
+It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction.
+Table 7 shows graph classification results.
+Training and inferring $G^{\prime}$ outperforms the baselines.
+Methods Models
+Reduction
+Ratio (r)
+Dataset
+ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
+Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
+GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
+SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
+GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
+FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
+GAT 0.1 $\mathbf{0.574}$ $\mathbf{0.705}$ $0.940$ $0.464$ $\mathbf{0.524}$
+SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
+GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
+Table 6: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
+Methods Models Graph per class Dataset
+AIDS PROTEINS
+DOSCOND GCN 1 $0.785$ $0.590$
+10 $0.666$ $0.647$
+50 $0.518$ $0.656$
+GAT 1 $0.608$ $0.652$
+10 $0.723$ $0.667$
+50 $0.729$ $0.657$
+KIDD GCN 1 $0.399$ $0.416$
+10 $0.236$ $0.659$
+50 $0.252$ $0.671$
+GAT 1 $0.441$ $0.416$
+10 $0.225$ $0.665$
+50 $0.218$ $0.671$
+SAGE 1 $0.653$ $0.416$
+10 $0.361$ $0.650$
+50 $0.283$ $0.673$
+GIN 1 $0.759$ $0.412$
+10 $0.637$ $0.655$
+50 $0.634$ $0.674$
+Methods Models
+Reduction
+Ratio (r)
+Dataset
+AIDS PROTEINS
+Full GCN 1.0 $0.788$ $0.710$
+GAT 1.0 $0.802$ $0.645$
+SAGE 1.0 $0.770$ $0.613$
+GIN 1.0 $0.800$ $0.774$
+FIT-GNN GCN 0.1 $0.810$ $0.783$
+0.3 $\mathbf{0.844}$ $\mathbf{0.826}$
+0.5 $0.836$ $0.696$
+0.7 $0.793$ $0.783$
+GAT 0.1 $0.784$ $0.652$
+0.3 $0.793$ $0.522$
+0.5 $0.810$ $0.739$
+0.7 $0.759$ $0.522$
+SAGE 0.1 $0.793$ $0.696$
+0.3 $0.828$ $0.696$
+0.5 $0.793$ $0.696$
+0.7 $0.819$ $\mathbf{0.826}$
+GIN 0.1 $0.819$ $0.652$
+0.3 $0.828$ $0.783$
+0.5 $0.793$ $0.478$
+0.7 $0.819$ $0.739$
+Table 7: Results on Graph Classification (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
+6.2 Inference Time and Memory
+As mentioned in Section 4.3, our inference time is less than the standard GNN if certain conditions are met.
+In Table 8(a) and 8(b), we club all baseline models into one, since all baselines are inferred on the whole graph.
+We empirically show the reduction of time and space by our FIT-GNN model during inference.
+For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
 FIT-GNN
-Dataset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
-ZINC (Subset) $0.688$ $\mathbf{0.578}$ $0.657$
-QM9 ($\mu$) $0.869$ $\mathbf{0.841}$ $0.948$
-QM9 ($\Delta\epsilon$) $1.012$ $\mathbf{0.875}$ $0.969$
-QM9 ($\mathrm{ZPVE}$) $1.091$ $\mathbf{0.818}$ $0.840$
-QM9 ($U^{\mathrm{ATOM}}$) $1.078$ $0.754$ $\mathbf{0.630}$
-Table 6: Graph regression tasks results on ZINC (Subset) and QM9.
-The metrics shown are in terms of normalized MAE (lower is better) with Gs-train-to-Gs-infer experimental setup.
-The Cluster Nodes method is used to append additional nodes to subgraphs.
-Gc-train-to
--Gs-infer
-Gs-train-to
--Gs-infer
-Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
-AIDS $0.792$ $0.810$ $0.836$ $0.812$ $\mathbf{0.843}$
-PROTEINS $0.623$ $0.636$ $0.681$ $0.812$ $\mathbf{0.821}$
-Table 7: Results for the graph classification task on AIDS and PROTEINS are presented.
-The metrics reflect accuracy; higher values indicate better performance.
-The table also compares two experimental setups.
-Cora Chameleon
-Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
-variation_neighborhoods $0.821\pm 0.003$ $0.803\pm 0.005$ $\mathbf{0.465\pm 0.005}$ $\mathbf{0.481\pm 0.006}$
-algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
-kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
-heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
-variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
-variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
-Table 8: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
-Table 3 shows the node regression error and standard deviation on coarsening ratios $0.1$ and $0.3$.
-All the experiments for node regression are shown with Cluster node method.
-With a higher coarsening ratio, more subgraphs are created.
-As the coarsening ratio increases, the size of each subgraph also reduces, and important neighboring node information gets lost.
-Hence, the performance decreases with an increase in the coarsening ratio.
-Table 4 shows the results for node classification accuracy and standard deviation on different coarsening ratios on the Cora dataset with Cluster Nodes approach.
-With coarsening ratios $0.1$ and $0.3$, the observed performance is better than other coarsening ratios, and the inference time taken is also less.
-This motivates us to experiment on other datasets with lower coarsening ratios.
-Table 5 shows the results for node classification on different datasets using coarsening ratios $0.1$ and $0.3$.
-From the results, it can be observed that the accuracy is comparable to the baseline.
-Since experimenting on the full OGBN-Products dataset is not feasible due to memory constraints, we create a proxy graph by taking the maximum-sized communities with the number of nodes that sum up to $165000$.
-For the subset of the OGBN-Products graph, the performance is better with the baseline even with $0.56\%$ training nodes.
-For larger datasets like Pubmed and Physics Coauthor, a lower coarsening ratio doesn’t perform better than the baseline.
-Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset.
-All the results show that the FIT-GNN model’s performance is comparable to the baselines.
-It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high.
-It implies that finer subgraphs lose out global information about the molecule, which is necessary for the prediction.
-Table 7 displays graph classification results using two experimental setups across two datasets.
-Training on coarsened graphs outperforms the baseline but discards important information.
-Conversely, training on subgraphs significantly improves predictions compared to the baseline.
-PROTEINS ZINC (subset)
-Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
-variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.629}$ $0.859$
-algebraic_JC $\mathbf{0.783}$ $0.478$ $0.823$ $0.810$
-kron $0.522$ $0.652$ $0.663$ $1.201$
-heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
-variation_edges $0.652$ $0.565$ $1.548$ $1.883$
-variation_cliques $0.696$ $0.652$ $0.689$ $\mathbf{0.742}$
-Table 9: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets.
-The metric for the PROTEINS dataset is accuracy (higher the better), and for the ZINC (subset) is normalized MAE (lower the better)
-Table 8 and Table 9 show the comparison of performance with different coarsening algorithms used by Huang et al.
-(2021).
-We find that variation_neighborhoods is more consistent among all other coarsening algorithms, therefore we use it to report other results also.
-5.3 Inference Time and Memory Comparison
-FIT-GNN
-Dataset Classical $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
+Dataset Baselines $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
 Chameleon $0.0027$ $0.0016$ $\mathbf{0.0014}$
@@
 OGBN-Products $0.1762$ $0.0017$ $\mathbf{0.0016}$
-Table 10: Inference time (sec) comparison for a single node using standard GNN as a baseline and FIT-GNN with two different coarsening ratios. Lower is better. Here, we used Cluster Nodes.
-Figure 3: Comparison of inference time memory consumption (in MB) for Coauthor Physics Dataset between baseline and FIT-GNN
+(a) Node-Level Tasks
 FIT-GNN
-Dastaset Classical $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
+Dastaset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
 ZINC (subset) $\mathbf{0.00184}$ $\mathbf{0.00184}$ $0.00190$
@@
 PROTEINS $0.00165$ $\mathbf{0.00160}$ $0.00163$
-Table 11: Inference time (sec) comparison for the graph-level task. We use the Cluster Nodes method to add additional nodes to each subgraph and Gc-train-to-Gc-infer experiment setup
-As mentioned in Section 4.3, our inference time is lesser than the standard GNN if certain conditions are met.
-Therefore, we empirically show reduction of time and space by our FIT-GNN model during inference.
-For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
-In Table 10, we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks.
+(b) Graph-Level Tasks
+Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks.
+We show single-node prediction time using two different coarsening ratios for Node-Level tasks.
+Here, we used Cluster Nodes for the node-level task.
+We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
+[…]
+Figure 3: GPU memory consumption (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference.
+In Table 8(a), we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks.
 The baseline model processes the entire graph, increasing inference time, especially for large graphs.
 In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions.
-For larger datasets like the OGBN-Products, we see up to $\mathbf{100\times}$ reduction in inference time.
-Table 17 in the Appendix shows the detailed comparison of memory consumptions for different datasets for both Extra Nodes and Cluster Nodes method along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{70\times}$ less memory than the baseline.
-Table 11 compares the inference time of graph classification and graph regression task.
-We randomly select $1000$ graphs from the test split and infer them for baseline and FIT-GNN.
-We also append additional nodes to each subgraph using Cluster Nodes method.
+For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity.
+Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges.
+We see up to $\mathbf{100\times}$.
+Section C elaborates more on the inference time for datasets with a larger number of nodes.
+Traditional methods fail to infer for very large datasets due to memory constraints.
+However, our approach can give an inference on the whole graph.
+Table 11 in the Appendix shows the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the baseline.
+Table 8(b) compares the inference time of graph classification and graph regression tasks.
+We predict for randomly selected $1000$ graphs from the test split.
 The table shows how our method is comparable and sometimes faster than the baseline.
@@
 Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges.
-In table 16 in the Appendix, we show the estimated $\phi_{\textup{max}}$ value from the feature dimension of the nodes and the coarsening ratio.
-Combinations with negative numbers indicate the FIT-GNN-based approach won’t reduce the inference time.
-Overall, the inference time and memory for all the tasks mentioned are drastically less than the classical approach while maintaining the performance mentioned in the previous section.
+Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
 7 Conclusion
@@
 We have provided theoretical insight corresponding to the number of nodes in the graph for which the FIT-GNN model reduces the time and space complexity.
-Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of the inference time and consuming a fraction of the memory.
-A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes.
-Appendix A Dataset Description
+Empirically, we have shown that our method is comparable to the uncoarsened baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory.
+A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes, and tackling real-world data-intensive challenges, for example, weather forecasting.
+Appendix
+Appendix A More on Extra Nodes and Cluster Nodes
+A.1 Extra Nodes
+Lemma 4.1. Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
+Proof.
+There are $2$ sets of nodes in $G_{i}$ as follows:
+  • The set $S_{1}$ of nodes with 1-hop neighbours in $G_{i}$.
+  • The set $S_{2}$ of nodes with not all 1-hop neighbors in $G_{i}$
+Let $\mathcal{I}^{1}_{i}$ be the number of nodes whose information does not get passed on after $1$ layer of GNN for $G_{i}$.
+$\mathcal{I}^{1}_{i}=\left|\bigcup_{v\in S_{2}}\mathcal{N}_{1}(v)-V(G_{i})\right|$
+[…]
+$\Rightarrow\left|\mathcal{E}_{G_{i}}\right|=\mathcal{I}^{1}_{i}$
+Hence, when Extra Nodes is used, 1 Layer GNN model cannot distinguish $G$ and $\mathcal{G}_{s}$. ∎
+To understand the information loss when taking $2$ Layers of GNN, we divide $G_{i}$ into $3$ sets of nodes.
+  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
+  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
+  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
+Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
+$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})\right|$
+When we use Extra Nodes, the information loss can be written as follows:
+$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})-\mathcal{E}_{G_{i}}\right|$
+The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other.
+An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Nodes method is used.
+A.2 Cluster Nodes
+Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$.
+Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$.
+This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
+Previously, according to Lemma 4.1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method.
+It was also easy to quantify in terms of the number of nodes.
+However, it is different for Cluster Nodes.
+Let us discuss the issues first.
+  • Only a weighted version of node information is shared with the subgraph.
+Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$.
+Then the information contributed by these nodes is $\frac{d_{a}x_{a}+d_{b}x_{b}}{\sum_{p}d_{p}}$.
+Here $d_{p}$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
+  • Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d_{p}x_{p}}{\sum_{p}d_{p}}$. This will capture further dependencies.
+The performance of Cluster Node will depend on some distance or similarity metric between $x_{c},x_{d}$ and $f(C_{j})$.
+Appendix B More Algorithms
+Algorithm 3 Training GNN (Train on $G^{\prime}$) [8]
+1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
+2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$;
+3:Construct Coarsened graph $G^{\prime}$ using $P$;
+4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{T}X$;
+5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{T}Y)$ ;
+[…]
+[…]
+8:Train $M$ to minimize LossLoss;
+Algorithm 4 Node-Model(L,A,D,X)(L,A,D,X)
+1:Number of Layers $L$; AA; $D$; $X^{(0)}=X$;
+2:for $i=1$ to $L$ do
+3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\quad\#$ Equation 1
+4:end for
+[…]
+6:return $Z$
+Algorithm 5 Graph-Model-$G^{\prime}(L,A^{\prime},D^{\prime},X^{\prime})$
+1:Number of Layers $L;A^{\prime};D^{\prime};X^{(0)}=X^{\prime}$
+2:for $i=1$ to $L$ do
+[…]
+4:end for
+[…]
+[…]
+7:return $Z$
+Appendix C More on Time and Space complexity
+Figure 4: The plot shows the feasibility region (green) corresponding to the coarsening ratio and Number of nodes (In order of Millions) in the graph for a fixed feature dimension of $d=200$.
+In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities.
+Figure 4 shows that those conditions can be satisfied for any dataset with nodes in order of millions.
+For node-level tasks, we show that the inference phase will take less time for single node prediction.
+To show that the time difference between the time taken by the classical approach and ours will increase as the number of nodes increases, we can formulate the time difference as follows:
+$\mathcal{T}(n)=\underset{\textup{Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}-\underset{\textup{ FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (3)
+In the above equation, we use the notations already defined in Section 4.3.
+We can also assume that $\max_{i}^{k}n_{i}=\frac{n}{\alpha}$, where $\alpha>1$ is some constant.
+Then we can rewrite the Equation 3 as:
+$\mathcal{T}(n)=n^{2}d+nd^{2}-(\frac{n}{\alpha}+\phi_{\max})^{2}d+(\frac{n}{\alpha}+\phi_{\max})d^{2}$
+We can observe that the above equation is strictly increasing for $n>0$. Hence the time difference also increases as the number of nodes in a dataset increases.
+Table 9: Upper Bound of $\phi_{\textup{max}}$ values for different datasets and coarsening ratios.
+Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
+PROTEINS $6.999$ $0.333$ 1.000-1.000 1.571-1.571
+AIDS $322.000$ $82.000$ $34.000$ $13.429$
+QM9 $79.000$ $19.000$ $7.000$ $1.857$
+ZINC (subset) 11.000-11.000 4.333-4.333 3.000-3.000 2.429-2.429
+Appendix D Dataset Description
 Dataset
@@
 Features Classes
-PROTEINS $1113$ $19$ $72$ $3$ $2$
-AIDS $2000$ $7$ $16$ $38$ $2$
-Table 12: Summary of datasets used for Graph classification Dataset
+PROTEINS $1113$ 1919 $72$ 33 22
+AIDS $2000$ 77 $16$ $38$ 22
+(a) Summary of datasets used for Graph classification Dataset
 Dataset
@@
 Targets
-QM9 $130831$ $8$ $18$ $11$ $19$
-ZINC (subset) $10000$ $11$ $25$ $1$ $1$
-Table 13: Summary of datasets used for Graph regression Dataset
+QM9 $130831$ 88 $18$ $11$ 1919
+ZINC (subset) $10000$ $11$ $25$ 11 11
+(b) Summary of datasets used for Graph regression Dataset
 Dataset Nodes Edges Features Classes
-Cora $2708$ $5278$ $1433$ $7$
-Citeseer $3327$ $4552$ $3703$ $6$
-Pubmed $19717$ $44324$ $500$ $3$
-DBLP $17716$ $52867$ $1639$ $4$
+Cora $2708$ $5278$ $1433$ 77
+Citeseer $3327$ $4552$ $3703$ 66
+Pubmed $19717$ $44324$ $500$ 33
+DBLP $17716$ $52867$ $1639$ 44
 Physics
 Coauthor
-$34493$ $247962$ $8415$ $5$
-OGBN-Products $165000$ $4340428$ $100$ $47$
-Table 14: Summary of datasets used for Node classification Dataset
+$34493$ $247962$ $8415$ 55
+OGBN-Products $2449029$ $61859140$ $100$ $47$
+(c) Summary of datasets used for Node classification Dataset
 Dataset Nodes Edges Features
@@
 Targets
-Chameleon $2277$ $31396$ $128$ $1$
-Squirrel $5201$ $198423$ $128$ $1$
-Crocodile $11631$ $170845$ $128$ $1$
-Table 15: Summary of datasets used for Node regression Dataset
-Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
-PROTEINS $6.999$ $0.333$ $-1.000$ $-1.571$
-AIDS $322.000$ $82.000$ $34.000$ $13.429$
-QM9 $79.000$ $19.000$ $7.000$ $1.857$
-ZINC (subset) $-11.000$ $-4.333$ $-3.000$ $-2.429$
-Table 16: Calculated $\phi_{\textup{max}}$ values for different datasets and coarsening ratios as per Lemma 1.
-Appendix B More on Extra Nodes and Cluster Nodes
-B.1 Extra Nodes
-There are $3$ sets of nodes in $G_{i}$.
-  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
-  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
-  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
-Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
-$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-V(G_{i})\right|$
-When we use Extra Nodes, the information loss can be written as follows:
-$\mathcal{I}^{2}_{i}=\left|\bigcup_{v\in S_{3}}\mathcal{N}_{2}(v)-\mathcal{E}_{G_{i}}\right|$
-The above entity will depend on the density of the subgraphs formed and the number of connections each subgraph shares with each other.
-An algorithm with an objective to reduce this entity for all subgraphs will lose the least amount of information when Extra Node method is used.
-B.2 Cluster Nodes
-Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$.
-Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$.
-This is one of the functions $f$ used to create the features of the cluster node from $C_{i}$.
-Previously, according to Lemma 1, there is no information loss when using the $1$ layer of GNN and Extra Nodes method.
-It was also easy to quantify in terms of the number of nodes.
-However, it is different for Cluster Nodes.
-Let us discuss the issues first.
-$\bullet$ Only a weighted version of node information is shared with the subgraph.
-Suppose $v_{c},v_{d}\in G_{i}$ is connected to $v_{a},v_{b}\in G_{j}$.
-Then the information contributed by these nodes is $\frac{d(v_{a})x_{a}+d(v_{b})x_{b}}{\sum_{p}d(v_{p})}$.
-Here $d(v_{p})$ represents the degree of node $v_{p}$, and $x_{p}$ represents the feature of node $v_{p}$.
-$\bullet$ Other node information will also be shared which is $\frac{\sum_{p\neq v_{a},v_{b}}d(v_{p})x_{p}}{\sum_{p}d(v_{p})}$. This will capture further dependencies.
-The performance of Cluster Node will depend on some distance or similarity metric between $x_{c},x_{d}$ and $f(C_{j})$.
-Appendix C Experiment Details (Parameters and Device Configuration)
+Chameleon $2277$ $31396$ $128$ 11
+Squirrel $5201$ $198423$ $128$ 11
+Crocodile $11631$ $170845$ $128$ 11
+(d) Summary of datasets used for Node regression Dataset
+Appendix E Experiment Details (Parameters and Device Configuration)
 For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight.
 We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks.
-For both coarsened graph and set of subgraph based model, we set epochs to 300 and the number of layers of GCN to 2 for training.
+For both coarsened graph (Gc-train) and set of subgraph based (Gs-train) model, we set epochs to 300 and the number of layers of GCN to 2 for training.
 We set the hidden dimensions for each layer of GCN to $512$.
-The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ $2.20$GHz, $256$GB RAM, NVIDIA A100 $40$GB GPU. We use Pytorch Geometric to to train our models.
-Appendix D Results
-This section shows the memory consumption for different datasets with different methods.
-Also, we show how the FIT-GNN model performs with different methods of appending additional nodes to subgraphs at different coarsening ratios.
-FIT-GNN
-Dataset Split $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$ Baseline
-Cora cluster $1.310$ $0.648$ $0.694$ $0.868$ $15.743$
-extra $2.072$ $1.134$ $1.017$ $1.017$
-Citeseer cluster $31.559$ $2.148$ $1.253$ $1.686$ $49.488$
-extra $31.559$ $3.114$ $1.788$ $1.788$
-Pubmed cluster $0.613$ $0.569$ $0.571$ $0.573$ $41.227$
-extra $0.798$ $0.590$ $0.590$ $0.590$
-DBLP cluster $6.067$ $2.940$ $1.847$ $1.966$ $118.121$
-extra $13.907$ $4.663$ $2.649$ $2.274$
-Physics Coauthor cluster $24.025$ $13.088$ $13.388$ $14.353$ $1169.521$
-extra $63.146$ $22.434$ $22.434$ $22.434$
-Chameleon cluster $0.212$ $0.246$ $0.291$ $0.438$ $2.198$
-extra $0.653$ $0.653$ $0.653$ $0.592$
-Crocodile cluster $1.183$ $1.188$ $1.735$ $1.866$ $11.562$
-extra $4.368$ $4.368$ $4.368$ $3.134$
-Squirrel cluster $1.507$ $1.607$ $1.903$ $3.009$ $9.075$
-extra $7.275$ $7.275$ $7.275$ $6.953$
-OGBN-Products cluster $6.047$ $5.613$ $8.636$ $208.029$
-Table 17: Summary of memory consumption of datasets used for node-level tasks. All units are in MegaBytes (MB)
+The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to to train our models.
+Appendix F Results
+Table 11 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes.
+Also, in Table 12 we show how the FIT-GNN model performs on Node Regression datasets for different coarsening ratios with Cluster Nodes method under Gs-train-to-Gs-infer setup.
+Dataset Appending Nodes FIT-GNN Baseline
+$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
+Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
+Extra Node $0.622$ $0.622$ $0.622$ $0.564$
+Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
+Extra Node $4.586$ $4.586$ $4.586$ $4.586$
+Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
+Extra Node $7.190$ $7.190$ $7.190$ $7.190$
+Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
+Extra Node $1.906$ $1.103$ $0.970$ $0.970$
+Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
+Extra Node $30.097$ $2.490$ $1.835$ $1.835$
+Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
+Extra Node $0.778$ $0.562$ $0.562$ $0.562$
+DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
+Extra Node $28.271$ $6.008$ $2.308$ $2.207$
+Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
+Extra Node $41.066$ $19.936$ $19.936$ $19.936$
+OGBN-Products Cluster Node $28.399$ $2840.706$
+Table 11: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)
+Methods Models Reduction Ratio (r) Dataset
+Chameleon Crocodile Squirel
+Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
+GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
+SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
+GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
+FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
+0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
+0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
+0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
+GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
+0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
+0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
+0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
+SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
+0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
+0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
+0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
+GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
+0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
+0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
+0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
+Table 12: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
+Methods Models Reduction Ratio (r) Dataset
+Cora Citeseer Pubmed DBLP Physics
+Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $0.933\pm 0.000$
+GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
+SGGC GCN 0.1 $0.752\pm 0.002$ $0.704\pm 0.003$ $0.730\pm 0.004$ $0.728\pm 0.019$ $0.892\pm 0.010$
+0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
+0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
+0.7 $0.805\pm 0.002$ $0.711\pm 0.003$ $0.786\pm 0.003$ $0.733\pm 0.014$ $0.931\pm 0.003$
+GAT 0.1 $0.630\pm 0.001$ $0.605\pm 0.025$ $89\pm 0.033 0.606$$06\pm 0.017$ $0.854\pm 0.008$
+0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
+0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
+0.7 $0.643\pm 0.019$ $0.570\pm 0.041$ $0.733\pm 0.023$ $0.613\pm 0.018$ $0.851\pm 0.026$
+GCOND GCN 0.1 $0.789\pm 0.003$ $0.676\pm 0.020$ $0.785\pm 0.003$ $0.718\pm 0.007$ $0.884\pm 0.008$
+0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
+0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
+0.7 $0.793\pm 0.000$ $0.680\pm 0.012$ $0.775\pm 0.002$ $0.751\pm 0.000$ $0.914\pm 0.004$
+GAT 0.1 $0.657\pm 0.067$ $0.553\pm 0.116$ $0.511\pm 0.079$ $0.475\pm 0.057$ OOM
+0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
+0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
+0.7 $0.706\pm 0.062$ $0.453\pm 0.118$ $0.416\pm 0.036$ $0.462\pm 0.033$ OOM
+BONSAI GCN 0.1 $0.729\pm 0.008$ $0.576\pm 0.006$ $0.600\pm 0.098$ $0.699\pm 0.031$ $0.822\pm 0.014$
+0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
+0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
+0.7 $0.705\pm 0.009$ $0.582\pm 0.007$ $0.474\pm 0.020$ $0.694\pm 0.023$ $0.844\pm 0.015$
+GAT 0.1 $0.685\pm 0.004$ $0.589\pm 0.002$ $0.446\pm 0.014$ $0.642\pm 0.015$ $0.803\pm 0.005$
+0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
+0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
+0.7 $0.703\pm 0.010$ $0.594\pm 0.002$ $0.497\pm 0.028$ $0.705\pm 0.047$ $0.799\pm 0.010$
+FIT-GNN GCN 0.1 $0.790\pm 0.004$ $0.701\pm 0.001$ $0.757\pm 0.002$ $0.755\pm 0.001$ $0.909\pm 0.000$
+0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $\mathbf{0.932\pm 0.000}$
+0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
+0.7 $0.813\pm 0.002$ $0.664\pm 0.001$ $0.756\pm 0.002$ $0.746\pm 0.001$ $0.925\pm 0.000$
+GAT 0.1 $0.773\pm 0.021$ $0.694\pm 0.003$ $0.737\pm 0.002$ $0.729\pm 0.010$ $0.887\pm 0.003$
+0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
+0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
+0.7 $0.797\pm 0.003$ $0.662\pm 0.003$ $0.759\pm 0.002$ $0.739\pm 0.002$ $0.903\pm 0.003$
+Table 13: Results for node classification tasks with accuracy as the metric (higher is better).
+We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
+Methods Models Reduction Ratio (r) Dataset
+ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
+Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
+GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
+SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
+GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
+FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
+0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
+0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
+0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
+GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $\mathbf{0.524}$
+0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
+0.5 $0.620$ $\mathbf{0.693}$ $1.227$ $0.475$ $0.561$
+0.7 $0.658$ - - - -
+SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
+0.3 $0.671$ - - - -
+0.5 $0.695$ - - - -
+0.7 $0.645$ - - - -
+GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
+0.3 $0.669$ - - - -
+0.5 $0.631$ - - - -
+0.7 $0.658$ - - - -
+Table 14: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
+Table 15: The Tables show the comparison between Extra Nodes and Cluster Nodes methods for node classification and node regression task.
+We show the average of the top 10 accuracies (higher the better) of the node classification task on Cora dataset and the average of the best 10 losses (lesser the better) of the node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup.
+They demonstrate the superior performance of the Cluster Nodes method over Extra Nodes.
+(a) Node Classification on Cora.
 r Extra Nodes Cluster Nodes
-$0.1$ $0.798\pm 0.008$ $\mathbf{0.824\pm 0.003}$
-$0.3$ $0.798\pm 0.008$ $\mathbf{0.807\pm 0.004}$
-$0.5$ $0.807\pm 0.004$ $\mathbf{0.814\pm 0.004}$
-$0.7$ $0.797\pm 0.007$ $\mathbf{0.808\pm 0.006}$
-Table 18: Average of top 10 accuracies (higher the better) of node classification task on Cora dataset with Gs-train-to-Gs-infer experimental setup.
-The table also demonstrates the superior performance of the Cluster Nodes method over Extra Nodes.
+$0.1$ $\mathbf{0.798\pm 0.008}$ $0.790\pm 0.004$
+$0.3$ $0.798\pm 0.008$ $\mathbf{0.800\pm 0.003}$
+$0.5$ $0.807\pm 0.004$ $\mathbf{0.829\pm 0.002}$
+$0.7$ $0.797\pm 0.007$ $\mathbf{0.813\pm 0.002}$
+(b) Node Regression on Chameleon.
 r Extra Nodes Cluster Nodes
-0.1 $0.504\pm 0.004$ $\mathbf{0.460\pm 0.003}$
-0.3 $0.579\pm 0.006$ $\mathbf{0.499\pm 0.006}$
-0.5 $0.585\pm 0.004$ $\mathbf{0.505\pm 0.004}$
-0.7 $\mathbf{0.504\pm 0.006}$ $0.560\pm 0.004$
-Table 19: Average of best 10 losses (lesser the better) of node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup.
-The table also demonstrates the better performance of the Cluster Nodes method over Extra Nodes.
-Dataset Coarsening Ratio Top 10 Loss
-Chameleon Baseline $0.844\pm 0.000$
-$0.1$ $\mathbf{0.460\pm 0.002}$
-$0.3$ $0.499\pm 0.006$
-$0.5$ $0.587\pm 0.000$
-$0.7$ $0.591\pm 0.080$
-Crocodile Baseline $0.853\pm 0.000$
-$0.1$ $\mathbf{0.362\pm 0.002}$
-$0.3$ $0.364\pm 0.001$
-$0.5$ $0.375\pm 0.001$
-$0.7$ $0.384\pm 0.002$
-Squirrel Baseline $0.809\pm 0.000$
-$0.1$ $\mathbf{0.648\pm 0.002}$
-$0.3$ $0.672\pm 0.003$
-$0.5$ $0.729\pm 0.004$
-$0.7$ $0.694\pm 0.004$
-Table 20: Node Regression Results with Gs-train-to-Gs-infer experimental setup.
+0.1 $\mathbf{0.504\pm 0.004}$ $0.506\pm 0.006$
+0.3 $0.579\pm 0.006$ $\mathbf{0.489\pm 0.003}$
+0.5 $0.585\pm 0.004$ $\mathbf{0.484\pm 0.004}$
+0.7 $\mathbf{0.504\pm 0.006}$ $0.539\pm 0.009$
+Table 16: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets.
+The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
+PROTEINS ZINC (subset)
+Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
+variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.573}$ $\mathbf{0.620}$
+algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
+kron $0.522$ $0.652$ $0.663$ $1.201$
+heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
+variation_edges $0.652$ $0.565$ $1.548$ $1.883$
+variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
+Table 17: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
+Cora Chameleon
+Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
+variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
+algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
+kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
+heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
+variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
+variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$

### v4

@@
 This scalability issue affects both the training and inference phases.
-Previous research has sought to mitigate the higher training computation by using three broad categories of methods: 1.
-Training every iteration on the sampled subgraph 2.
+Previous research has sought to mitigate the high training computation cost by using three broad categories of methods: 1.
+Training every iteration on a sampled subgraph 2.
 Training on a coarsened graph generated from a coarsening algorithm 3.
@@
 The aim of reducing computational costs has been approached through various methods.
-Research by Chiang et al.
-[1] and Zeng et al.
-[2] uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training.
-Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen and Zhu [3], Chen et al.
-[4], Cong et al.
-[5], Zou et al.
-[6].
-Additionally, Fahrbach et al.
-[7] introduced a coarsening algorithm based on Schur complements to create embeddings for vertices in large graphs, while Huang et al.
-[8] (SGGC) applied GNNs to coarsened graphs, reducing nodes and edges to lower training time.
-Kumar et al.
-[9] proposed feature-based coarsening algorithms.
-Jin et al.
-[10] (GCOND), Wang et al.
-[11] (GC-SNTK) presented algorithms to generate a synthetic graph that mimics the original graph.
+Research by 10.1145/3292500.3330925 and Zeng2020GraphSAINT: uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training.
+Other methods, including layer-wise sampling and mini-batch training, have been explored by chen2018stochastic; chen2018fastgcn; 10.1145/3394486.3403192; DBLP:journals/corr/abs-1911-07323.
+Additionally, DBLP:journals/corr/abs-2007-02817 introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while 10.1145/3447548.3467256 [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time.
+pmlr-v202-kumar23a; joly2024graph; NEURIPS2024_733209a1; 10.1145/3589335.3651920 also proposed different coarsening algorithms for scalable training approaches.
+jin2021graph [GCOND], wang2024fast [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph.
 However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden.
-Gupta et al.
-[12] (BONSAI) addressed this issue by looking into computational trees and training on just the relevant computational trees.
+gupta2025bonsai [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees.
 However, these approaches do not decrease inference time since the inference phase takes the whole graph as input.
-Later, Xue et al.
-[13] shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach.
-In the graph-level task, Jin et al.
-[14] (DOSCOND) and Xu et al.
-[15] (KIDD) tackled the graph classification problem following up on the approach of GCOND and GC-SNTK.
+DBLP:journals/corr/abs-2202-00075 shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach.
+In the graph-level task, Jin_2022 [DOSCOND] and 10.1145/3580305.3599398 [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK.
 However, the graph regression problem was not explored in this direction.
@@
 The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors.
-Kipf and Welling [16] introduced Graph Convolutional Network (GCN) as follows:
+kipf2017semisupervised introduced Graph Convolutional Network (GCN) as follows:
 $X^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(l)}\mathcal{W}^{(l)})$ (1)
@@
 3.3 Graph Coarsening and Graph Partitioning
-Loukas [17] discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
+loukas2018graphreductionspectralcut discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
+$X^{\prime}$ is the coarsened node representation.
 We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes.
-Given a coarsening ratio $r\in[0,1]$, we have $k=n\times r$.
+Given a coarsening ratio $r\in[0,1]$, we have $n\times r, kk\in\mathbb{Z}$.
 We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes.
@@
 Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$.
-SGGC used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
+SGGC (10.1145/3447548.3467256) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
 Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$.
@@
 images/img2
-Figure 2: Methods of appending additional nodes in $G_{1},G_{2},G_{3}$.
+Figure 2: Figure showing the comparison between the Extra Node Method and the Cluster Node Method of appending additional nodes in $G_{1},G_{2},G_{3}$.
 4 Methodology
@@
   • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
-Xue et al.
-[13] introduced this approach to reduce information loss after partitioning.
-We add an edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
+DBLP:journals/corr/abs-2202-00075 introduced this approach to reduce information loss after partitioning.
+We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
+Node embedding of $v\in\mathcal{E}_{G_{i}}$ is $x_{v}\in X$.
+    $\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}$ (2)
   • Cluster Nodes: Although adding $1$-hop neighboring nodes in the subgraph helps recover lost information after partition, for multi-layered GNN models, information loss will persist.
 Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$.
-Liu et al.
-[18] introduced this method to add nodes in the graph instead of adding to separate subgraphs.
+liu2024scalable introduced this method to add nodes in the graph instead of adding to separate subgraphs.
 They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph.
-In our work, we add cross-cluster edges but do not sample from the constructed set of subgraphs.
-$\mathcal{E}_{G_{i}}=\bigcup_{v\in G_{i}}\left\{u\in V:u\in\mathcal{N}_{1}(v)\wedge u\notin G_{i}\right\}\quad\text{and}\quad\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (2)
+In our work, we add cross-cluster edges.
+For each node $t\in\mathcal{C}_{G_{i}}$, the node embedding $x_{t}=X^{\prime}_{t}$.
+    $\mathcal{C}_{G_{i}}=\bigcup_{v\in\mathcal{E}_{G_{i}}}\left\{t:P_{v,t}\neq 0\right\}$ (3)
 Figure 2 illustrates appending these additional nodes using these approaches.
@@
 Lemma 4.1.
-Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used.
-Proof.
-In Appendix A.1. ∎
+Models with $1$ layer of GNN cannot distinguish between $G$ and $\mathcal{G}_{s}$ when Extra Nodes method is used. (Proof in Appendix A.1)
 Using the Cluster Nodes approach to append additional nodes to subgraphs is better for two reasons:
-  • From Equation 2, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$ This is because, using Extra Nodes, we add all the neighboring nodes.
+  • From Equation 2 and 3, we can say that $\sum_{i}|\mathcal{E}_{G_{i}}|\geq\sum_{i}|\mathcal{C}_{G_{i}}|$.
+This is because, using Extra Nodes, we add all the neighboring nodes.
 However, using Cluster Nodes, we only add one representative node for each neighboring cluster.
@@
 4.3 Time and Space Complexity
+Table 1: Training and inference time and space complexity of FIT-GNN compared to classical and SGGC approaches.
 (a) Time Complexity
@@
 FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $d^{2}+\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]$
-Table 1: Training and inference time and space complexity of FIT-GNN compared to classical and SGGC approaches.
-The dimensions of the matrices in Equation 1 are $(n\times n)$ for $\tilde{A}$, $(n\times d)$ for $X^{(l)}$ and $(d\times d)$ for $\mathcal{W}^{(l)}$.
+In Equation 1, the matrix dimensions are as follows: $\tilde{A}$ has dimensions $(n\times n)$, $X^{(l)}$ is of dimensions $(n\times d)$, and $\mathcal{W}^{(l)}$ has dimensions $(d\times d)$.
 Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$.
@@
 The complexity is based on GCN.
-For other architectures like GAT [19], the complexity changes.
+For other architectures like GAT veličković2018graphattentionnetworks, the complexity changes.
 SGGC improved the time and space complexity (Table 1) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
@@
 One is the classical model, where no coarsening or partitioning is done; the second is SGGC, where training is performed on a smaller graph $G^{\prime}$ and inference on $G$; the third is our approach, FIT-GNN, where both training and inference are performed on $\mathcal{G}_{s}$.
-The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{\textup{max}}$ where $\phi_{\textup{max}}=\max_{i}(\mathcal{E}_{G_{i}})$ for Extra Nodes or $\phi_{\textup{max}}=\max_{i}(\mathcal{C}_{G_{i}})$ for Cluster Nodes is the maximum number of additional nodes added per subgraph.
+The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{i}$ and $\phi_{i}$ is the number of additional nodes appended in each subgraph.
 Table 1 compares different approaches.
 Lemma 4.2.
-The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$ for $r\leq\frac{d-2}{d+\phi_{\textup{max}}}$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}.$
+The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$ if $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}$.
 Proof.
-$\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{\textup{max}}^{2}d+2n_{i}\phi_{\textup{max}}d+n_{i}d^{2}+\phi_{\textup{max}}d^{2}]$
-$=k\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+k\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d$
-$=nr\phi_{\textup{max}}^{2}d+2n\phi_{\textup{max}}d+nd^{2}+nr\phi_{\textup{max}}d^{2}+\sum_{i=1}^{k}n_{i}^{2}d\quad(\textup{where }k=nr)$
-$=nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d$
-Now we know, $r\leq\frac{d-2}{d+\phi_{\textup{max}}}\Rightarrow r\phi_{\textup{max}}+2+rd\leq d$ and $\phi_{\textup{max}}\leq\frac{n^{2}-\sum_{i=1}^{k}n_{i}^{2}}{nd}\Rightarrow n\phi_{\textup{max}}d\leq n^{2}-\sum_{i=1}^{k}n_{i}^{2}$.
-Using this, we can say,
-$nd^{2}+n\phi_{\textup{max}}d(r\phi_{\textup{max}}+2+rd)+\sum_{i=1}^{k}n_{i}^{2}d\leq nd^{2}+(n^{2}-\sum_{i=1}^{k}n_{i}^{2})d+\sum_{i=1}^{k}n_{i}^{2}d=nd^{2}+n^{2}d\\$
-Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{\textup{max}})^{2}d+(n_{i}+\phi_{\textup{max}})d^{2}]\leq n^{2}d+nd^{2}$. ∎
+$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{i}^{2}d+2n_{i}\phi_{i}d+n_{i}d^{2}+\phi_{i}d^{2}]$
+$=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
+Now, define $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$.
+We also define the variances and covariance,
+$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2}$ $\qquad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},$
+$\text{Cov}(n_{i},\phi_{i})=\text{Cov}$ $:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
+We can also write,
+$\text{Var}(n_{i}+\phi_{i})=\frac{1}{k}\sum_{i=1}^{k}(n_{i}+\phi_{i})^{2}-(\mathbb{E}[n_{i}]+\mathbb{E}[\phi_{i}])^{2}=\text{Var}(n_{i})+\text{Var}(\phi_{i})+2\text{Cov}(n_{i},\phi_{i})$
+Using the definitions above, we can say:
+$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
+$=nd^{2}+d\left[k(\textup{Var}(n_{i})+\mathbb{E}[n_{i}]^{2})+k(\textup{Var}(\phi_{i})+\mathbb{E}[\phi_{i}]^{2})+2k(\textup{Cov}(n_{i},\phi_{i})+\mathbb{E}[n_{i}]\mathbb{E}[\phi_{i}])+dk\mathbb{E}[\phi_{i}]\right]$
+$=nd^{2}+nd\left[r\textup{Var}(n_{i})+\frac{1}{r}+r\textup{Var}(\phi_{i})+r\mathbb{E}[\phi_{i}]^{2}+2r\textup{Cov}(n_{i},\phi_{i})+2\mathbb{E}[\phi_{i}]+rd\mathbb{E}[\phi_{i}]\right]$
+$=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
+Given, $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}\Rightarrow\mathbb{E}[\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-(\frac{d}{2}+\frac{1}{r})$.
+Using this, we can say the following:
+$\sum_{i=1}^{k}[(n_{i}+$ $\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+nd\left[r\mathbb{E}[\phi_{i}]^{2}+(2+rd)\mathbb{E}[\phi_{i}]+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
+$\leq nd^{2}+nd\left[r\left(\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})+\frac{d^{2}}{4}+\frac{1}{r^{2}}+\frac{d}{r}-2\left(\frac{d}{2}+\frac{1}{r}\right)\sqrt{\Delta}\right)\right$.
+$\qquad\qquad\left.+2\sqrt{\Delta}-2\left(\frac{d}{2}+\frac{1}{r}\right)+rd\sqrt{\Delta}-rd\left(\frac{d}{2}+\frac{1}{r}\right)+r\textup{Var}(n_{i}+\phi_{i})+\frac{1}{r}\right]$
+$\qquad\qquad\qquad(\textup{where }\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})=\Delta)$
+$=nd^{2}+n^{2}d$
+Hence, $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$. ∎
+Corollary 4.3.
+$\mathbb{E}[\phi_{i}]$ has a positive upper bound when $\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$.
+Proof.
+From Lemma 4.2, we know the upper bound of $\mathbb{E}[\phi_{i}]$ is $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)$.
+If $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$, then we can say the following:
+$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$
+$\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$
+$\Rightarrow\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$
 For a given graph, when the above conditions are satisfied, our approach has better inference time and space complexity than other approaches.
-In Section C, we show that, for datasets with millions of nodes, the conditions outlined in Lemma 4.2 can indeed be satisfied.
-As the number of nodes increases, the gap in inference time between baseline methods and FIT-GNN widens.
-In table 9, we show the upper bound of $\phi_{\max}$ calculated as per the conditions mentioned in Lemma 4.2 for different datasets corresponding to different coarsening ratios.
-Negative upper bound is observed for some datasets, implying that our method won’t give better computational complexities in those scenarios.
+From Corollary 4.3, we can understand that it is ideal to have similarly sized subgraphs.
+For more details on time and space complexity, refer to Section C in the Appendix.
 5 Experiments
@@
   • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$.
+Our aim was to examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
   • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights.
+We explore whether training solely on the $G^{\prime}$ is sufficient for inference on $\mathcal{G}_{s}$.
+This setting presents a computationally efficient option, allowing users with limited resources to obtain results close to the best reported.
   • Gs-train-to-Gs-infer: Perform subgraph-level training and inference on $\mathcal{G}_{s}$.
-  • Gc-train-to-Gc-infer: Unlike node-level tasks, we can infer on $G^{\prime}$ (Gc-infer) itself for graph-level tasks after training on $G^{\prime}$.
+  • Gc-train-to-Gc-infer: Unlike node-level tasks, for graph-level tasks, inference can be done directly on $G^{\prime}$ because the label corresponds to the entire graph and not to individual nodes.
+Therefore, even with a reduced graph representation, meaningful inference is possible.
 Therefore, in this setup for graph-level tasks only, we train and infer on $G^{\prime}$.
@@
 For the node regression task, we only perform Gs-train-to-Gs-infer because a coarsened graph is not created for node regression datasets.
-[…]
-Tasks Datasets
-Node Regression
-(Train=30%, Val=20%, Test=50%)
-Chameleon [20]
-Crocodile [20]
-Squirrel [20]
-Node Classification
-(Train=20*c, Val=30*c, Test=Rest)
-Cora [21]
-Citeseer [22]
-Pubmed [23]
-Coauthor Physics [24]
-DBLP [25]
-OGBN-Products [26]
-Graph Regression
-(Train=50%, Val=25%, Test=25%)
-QM9 [27]
-ZINC (Subset) [28]
-Graph Classification
-(Train=50%, Val=25%, Test=25%)
-PROTEINS [29]
-AIDS [29]
-Table 2: The table shows the datasets used and the splits in each dataset.
-‘c’ represents the number of classes for the dataset.
-We applied the public split “fixed” from [30] for Cora, Citeseer, and PubMed.
-For the “random” split in node classification, we keep 2020 nodes per class for training, $30$ nodes for validation, and the rest as test nodes.
-[31] grouped the 1919 properties of each molecule in QM9 dataset into four broad categories.
-We predict one property from each broad category: the dipole moment ($\mu$), the gap between $\epsilon_{\mathrm{HOMO}}$ and $\epsilon_{\mathrm{LUMO}}$, zero-point vibration energy ($\mathrm{ZPVE}$), and at $298.15\mathrm{K}$, the atomization energy ($U^{\mathrm{ATOM}}$) of each molecule.
+Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: FIT-GNN.
+Table 2: Datasets and corresponding splits used in our experiments.
+‘c’ denotes the number of classes in node classification datasets.
+Public split “fixed” from 10.5555/3045390.3045396 is used for Cora, Citeseer, and PubMed.
+For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing.
+For QM9, we follow DBLP:journals/corr/GilmerSRVD17 and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
+Task Dataset (Reference) Train Val Test
+Node Regression Chameleon rozemberczki2019multiscale 30% 20% 50%
+Crocodile rozemberczki2019multiscale 30% 20% 50%
+Squirrel rozemberczki2019multiscale 30% 20% 50%
+Node Classification Cora McCallum2000 $20\times c$ $30\times c$ Rest
+Citeseer 10.1145/276675.276685 $20\times c$ $30\times c$ Rest
+Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 $20\times c$ $30\times c$ Rest
+Coauthor Physics DBLP:journals/corr/abs-1811-05868 $20\times c$ $30\times c$ Rest
+DBLP 10.1145/1401890.1402008 $20\times c$ $30\times c$ Rest
+OGBN-Products 10.5555/3495724.3497579 $20\times c$ $30\times c$ Rest
+Graph Regression QM9 DBLP:journals/corr/WuRFGGPLP17 50% 25% 25%
+ZINC (Subset) DBLP:journals/corr/Gomez-Bombarelli16 50% 25% 25%
+Graph Classification PROTEINS Morris+2020 50% 25% 25%
+AIDS Morris+2020 50% 25% 25%
 6 Result and Analysis
-Methods Models
-Reduction
-Ratio (r)
-Dataset
+6.1 Model Performance
+Table 3: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
+variation_neighborhoods coarsening algorithm is used.
+Methods Models Reduction Ratio (r) Dataset
 Chameleon Crocodile Squirel
@@
 GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
-FIT-GNN GCN 0.1 $\mathbf{0.496\pm 0.005}$ $0.364\pm 0.001$ $0.663\pm 0.003$
+FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
+0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
+0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
+0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
 GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
+0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
+0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
+0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
 SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
+0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
+0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
+0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
 GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
-Table 3: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
+0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
+0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
+0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
+Table 4: Results for node classification tasks with accuracy as the metric (higher is better).
+We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
 variation_neighborhoods coarsening algorithm is used.
@@
 0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
-Table 4: Results for node classification tasks with accuracy as the metric (higher is better).
-We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
-variation_neighborhoods coarsening algorithm is used.
-6.1 Model Performance
-Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$
+Node Regression: Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE).
+As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost.
+Hence, an increase in the coarsening ratio decreases performance.
+Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$. variation_neighborhoods coarsening algorithm is used. OOM means Out Of Memory.
 Full $0.873\pm 0.000$
@@
 FIT-GNN $\mathbf{0.894\pm 0.000}$
-Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE) with coarsening ratios $0.1$.
-Results using different coarsening ratios are in Table 12.
-As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost.
-Hence, an increase in the coarsening ratio decreases performance.
-Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach.
+Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach.
 Comparison with other coarsening ratios is in Table 10.
 Results show that the accuracy is comparable to the baselines.
+Figure 5 shows the comparison of performance between different experimental setups for different methods of appending nodes on the Cora dataset.
+From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task.
+For the ’None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance.
+We also observe in the ’None’ method that, with an increase in the coarsening ratio, the performance drops.
+This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss.
+For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases.
+This can be argued because of the change in the distribution of the coarsening graph and the set of subgraphs.
+When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance.
+While comparing the methods of appending nodes, we infer that the Cluster Node method is performing similarly compared to the Extra Nodes method, if not better.
 Table 5 shows results on the OGBN-Products dataset.
@@
 For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
-Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset.
-More results at Table 14.
-All the results show that the FIT-GNN model’s performance is better than the baselines.
-It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is high.
-It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction.
-Table 7 shows graph classification results.
-Training and inferring $G^{\prime}$ outperforms the baselines.
-Methods Models
-Reduction
-Ratio (r)
-Dataset
+Table 6: Results on Graph Regression with MAE as metric (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
+Methods Models Reduction Ratio (r) Dataset
 ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
@@
 FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
-GAT 0.1 $\mathbf{0.574}$ $\mathbf{0.705}$ $0.940$ $0.464$ $\mathbf{0.524}$
+0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
+0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
+0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
+GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $0.524$
+0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
+0.5 $0.620$ $0.693$ $1.227$ $0.475$ $0.561$
+0.7 $0.658$ $\mathbf{0.686}$ $1.030$ $0.491$ $0.703$
 SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
+0.3 $0.671$ $0.748$ $0.711$ $0.602$ $\mathbf{0.518}$
+0.5 $0.695$ $0.739$ $0.747$ $0.606$ $0.833$
+0.7 $0.645$ $0.722$ $0.677$ $0.659$ $1.005$
 GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
-Table 6: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
+0.3 $0.669$ $0.723$ $0.761$ $0.509$ $0.854$
+0.5 $0.631$ $0.725$ $0.848$ $0.737$ $1.058$
+Graph Regression: Table 6 shows the results for the graph regression task on ZINC (Subset) and QM9 dataset.
+All the results show that the FIT-GNN model’s performance is better than the baselines.
+It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large.
+It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction.
+Similarly, when inferred on a coarsened graph for each graph in the dataset, the performance degrades.
+This is because, when a graph is reduced to a smaller graph, the positional information is lost, which is crucial since in graphs representing different molecules, the position of a certain component in a molecule results in different properties of the molecule.
+Table 7: Results on Graph Classification with accuracy as metric (higher is better).
+We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup.
+algebraic_JC coarsening algorithm is used.
 Methods Models Graph per class Dataset
@@
 0.7 $0.819$ $0.739$
-Table 7: Results on Graph Classification (higher is better). We use the extra nodes method and Gc-train-to-Gc-infer as experimental setup. algebraic_JC coarsening algorithm is used.
+Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$.
+For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data.
+DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution.
+Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class.
+However, the graph size is kept similar to the graph size in the training set.
+These approaches also don’t cover the test set.
+Compared to these baselines, our approach reduces each graph into a set of subgraphs and a coarsened graph based on a reduction ratio.
+As a result, each graph in the training set and test set gets reduced.
+From our results, we see that FIT-GNN outperforms all the baselines.
+Table 11 and 12 present the ablation study on various coarsening algorithms for all the aforementioned tasks.
+From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others.
+Therefore, we report results on all datasets using variation_neighborhoods coarsening algorithm.
 6.2 Inference Time and Memory
@@
 For recording the inference time, we use the Python time package to calculate the difference in time before and after the inference step.
+Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks.
+We show single-node prediction time using two different coarsening ratios for Node-Level tasks.
+Here, we used Cluster Nodes for the node-level task.
+We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
 FIT-GNN
@@
 (b) Graph-Level Tasks
-Table 8: Inference time (sec) (Lower is better) comparison for Node-Level tasks and Graph-Level tasks.
-We show single-node prediction time using two different coarsening ratios for Node-Level tasks.
-Here, we used Cluster Nodes for the node-level task.
-We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
 […]
-Figure 3: GPU memory consumption (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference.
-In Table 8(a), we show the average time to predict for $1000$ nodes in the baseline and FIT-GNN models for node regression and classification tasks.
+Figure 3: GPU memory consumption in MegaBytes (MB) (log scale) for FIT-GNN (Cluster Node) at different reduction ratios $r$ and Baseline, during inference.
+variation_neighborhoods coarsening algorithm is used.
+In Table 8(a), we show the average time to predict for $1000$ queries (Prediction for Node $v_{i}$) in the Baseline and FIT-GNN models for node regression and classification tasks.
 The baseline model processes the entire graph, increasing inference time, especially for large graphs.
@@
 However, our approach can give an inference on the whole graph.
-Table 11 in the Appendix shows the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the baseline.
+Figure 3 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline.
+Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset.
+As the coarsening ratio increases, the time also increases, since the number of subgraphs increases with the coarsening ratio.
 Table 8(b) compares the inference time of graph classification and graph regression tasks.
@@
 Also, we observe that the inference time increases with a higher coarsening ratio.
-Because, with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges.
+Because with a higher coarsening ratio, the number of nodes in $G^{\prime}$ increases, resulting in more edges.
 Overall, the inference time and memory for all the tasks mentioned are drastically less than the baselines while maintaining the performance.
@@
 Appendix B More Algorithms
-Algorithm 3 Training GNN (Train on $G^{\prime}$) [8]
+Algorithm 3 Training GNN (Train on $G^{\prime}$) (10.1145/3447548.3467256)
 1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
@@
 Appendix C More on Time and Space complexity
-Figure 4: The plot shows the feasibility region (green) corresponding to the coarsening ratio and Number of nodes (In order of Millions) in the graph for a fixed feature dimension of $d=200$.
 In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities.
-Figure 4 shows that those conditions can be satisfied for any dataset with nodes in order of millions.
-For node-level tasks, we show that the inference phase will take less time for single node prediction.
-To show that the time difference between the time taken by the classical approach and ours will increase as the number of nodes increases, we can formulate the time difference as follows:
-$\mathcal{T}(n)=\underset{\textup{Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}-\underset{\textup{ FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (3)
-In the above equation, we use the notations already defined in Section 4.3.
-We can also assume that $\max_{i}^{k}n_{i}=\frac{n}{\alpha}$, where $\alpha>1$ is some constant.
-Then we can rewrite the Equation 3 as:
-$\mathcal{T}(n)=n^{2}d+nd^{2}-(\frac{n}{\alpha}+\phi_{\max})^{2}d+(\frac{n}{\alpha}+\phi_{\max})d^{2}$
-We can observe that the above equation is strictly increasing for $n>0$. Hence the time difference also increases as the number of nodes in a dataset increases.
-Table 9: Upper Bound of $\phi_{\textup{max}}$ values for different datasets and coarsening ratios.
-Dataset $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.7}$
-PROTEINS $6.999$ $0.333$ 1.000-1.000 1.571-1.571
-AIDS $322.000$ $82.000$ $34.000$ $13.429$
-QM9 $79.000$ $19.000$ $7.000$ $1.857$
-ZINC (subset) 11.000-11.000 4.333-4.333 3.000-3.000 2.429-2.429
+For node-level tasks, to obtain lower FIT-GNN inference time complexity as compared to classical GNN, we want the following as per asymptotic analysis,
+  1. 1.
+    Baseline vs. Single-Node Inference:
+    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]}}$ (4)
+  2. 2.
+    Baseline vs. Full-Graph Inference:
+    $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\sum_{i=1}^{k}\bigl[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}\bigr]}}$ (5)
+Figure 4: The figure shows the feasibility of our inference methods for different node-level datasets against coarsening ratios.
+Two important inference setups for each dataset: single-node inference with the computational cost $\mathcal{O}\left(\max_{i}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})+n\right)$, and full-graph inference with computational cost $\mathcal{O}\left(\sum_{i=1}^{k}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})\right)$, are compared against the baseline (classical GNN) computational cost of $\mathcal{O}\left(n^{2}d+nd^{2}\right)$.
+The coarsening method used is variation_neighborhoods.
+Both axes are in logarithmic scale (base 10).
+Figure 4 demonstrates the feasibility of our FIT-GNN method for different coarsening ratios with regards to full-graph inference time complexity.
+We empirically compute the LHS (baseline is classical GNN) and RHS of the Inequality 4 and 5for various coarsening ratios and plot the results for multiple datasets.
+The coarsening algorithm used is variation_neighborhoods.
+We make the following inferences:
+  1. 1.
+    There is a clear dataset-dependent computational-cost trade-off across different coarsening ratios.
+Moreover, the minimum inference time for our approach occurs at different coarsening ratios for different datasets.
+  2. 2.
+    Under single-node inference setup, FIT-GNN demonstrates increasingly superior inference time performance as compared to classical GNN with increasing coarsening ratio.
+This is because the individual subgraph sizes decrease with increasing coarsening ratio.
+  3. 3.
+    We observe lower inference time complexity of Cluster_node method as compared to Extra_node method under both inference setups.
+This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.
 Appendix D Dataset Description
@@
 Appendix F Results
+Figure 5 presents the ablation study conducted to compare the performance of FIT-GNN for different experimental setup using different methods for appending nodes in the subgraph.
+The experiments are conducted on the Cora dataset for different coarsening ratios.
+Figure 5: The plot shows the ablation study conducted on the Cora dataset to determine which experimental setup performs better than the others.
+The plot also compares different methods of appending nodes to the subgraphs and how performance changes with varying coarsening ratios.
+variation_neighborhoods coarsening algorithm is used.
+Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios.
+It is observed that the None method takes the least amount of time, which is obvious since the subgraphs are not appended with additional nodes.
+For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended.
+For Cluster Node, we observe the maximum time taken since not only are we looking at the neighboring nodes that are lost due to partition, but also we are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
+Figure 6: The plot shows how the coarsening time varies for the Cora dataset for different coarsening ratios using different methods of appending nodes to the subgraphs.
+variation_neighborhoods coarsening algorithm is used.
+Table 10 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$.
+To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 11 and 12.
 Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes.
-Also, in Table 12 we show how the FIT-GNN model performs on Node Regression datasets for different coarsening ratios with Cluster Nodes method under Gs-train-to-Gs-infer setup.
-Dataset Appending Nodes FIT-GNN Baseline
-$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
-Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
-Extra Node $0.622$ $0.622$ $0.622$ $0.564$
-Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
-Extra Node $4.586$ $4.586$ $4.586$ $4.586$
-Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
-Extra Node $7.190$ $7.190$ $7.190$ $7.190$
-Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
-Extra Node $1.906$ $1.103$ $0.970$ $0.970$
-Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
-Extra Node $30.097$ $2.490$ $1.835$ $1.835$
-Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
-Extra Node $0.778$ $0.562$ $0.562$ $0.562$
-DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
-Extra Node $28.271$ $6.008$ $2.308$ $2.207$
-Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
-Extra Node $41.066$ $19.936$ $19.936$ $19.936$
-OGBN-Products Cluster Node $28.399$ $2840.706$
-Table 11: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)
-Methods Models Reduction Ratio (r) Dataset
-Chameleon Crocodile Squirel
-Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
-GAT 1.0 $0.846\pm 0.002$ $0.850\pm 0.000$ $0.805\pm 0.001$
-SAGE 1.0 $0.848\pm 0.002$ $0.850\pm 0.000$ $0.804\pm 0.000$
-GIN 1.0 $0.843\pm 0.000$ $0.852\pm 0.000$ $0.796\pm 0.000$
-FIT-GNN GCN 0.1 $0.496\pm 0.005$ $0.364\pm 0.001$ $0.663\pm 0.003$
-0.3 $0.531\pm 0.006$ $0.369\pm 0.002$ $0.682\pm 0.002$
-0.5 $0.531\pm 0.004$ $0.371\pm 0.001$ $0.722\pm 0.002$
-0.7 $0.536\pm 0.009$ $0.368\pm 0.003$ $0.696\pm 0.002$
-GAT 0.1 $0.582\pm 0.029$ $0.395\pm 0.022$ $0.734\pm 0.019$
-0.3 $0.537\pm 0.021$ $0.383\pm 0.020$ $0.724\pm 0.020$
-0.5 $0.556\pm 0.018$ $0.386\pm 0.018$ $0.728\pm 0.023$
-0.7 $0.576\pm 0.018$ $0.395\pm 0.012$ $0.746\pm 0.020$
-SAGE 0.1 $0.506\pm 0.006$ $\mathbf{0.309\pm 0.014}$ $\mathbf{0.589\pm 0.003}$
-0.3 $0.489\pm 0.003$ $0.315\pm 0.001$ $0.602\pm 0.003$
-0.5 $\mathbf{0.484\pm 0.004}$ $0.317\pm 0.001$ $0.618\pm 0.002$
-0.7 $0.539\pm 0.009$ $0.323\pm 0.000$ $0.621\pm 0.004$
-GIN 0.1 $0.769\pm 0.101$ $0.722\pm 0.098$ $0.795\pm 0.018$
-0.3 $0.808\pm 0.056$ $0.697\pm 0.122$ $0.796\pm 0.004$
-0.5 $0.810\pm 0.062$ $0.730\pm 0.091$ $0.786\pm 0.032$
-0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
-Table 12: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
+Note that this memory consumption is to store the graph and the weight matrix in the memory only.
 Methods Models Reduction Ratio (r) Dataset
@@
 We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
-Methods Models Reduction Ratio (r) Dataset
-ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
-Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
-GAT 1.0 $0.736$ $0.919$ $1.015$ $1.104$ $1.066$
-SAGE 1.0 $0.685$ $0.885$ $1.015$ $1.109$ $1.076$
-GIN 1.0 $0.748$ $0.925$ $1.053$ $1.126$ $1.078$
-FIT-GNN GCN 0.1 $0.625$ $0.713$ $1.223$ $\mathbf{0.447}$ $1.223$
-0.3 0.573 $0.708$ $0.975$ $0.473$ $1.220$
-0.5 $0.645$ $0.702$ $1.153$ $0.485$ $1.179$
-0.7 $0.651$ $0.715$ $1.189$ $0.456$ $1.240$
-GAT 0.1 $0.574$ $0.705$ $0.940$ $0.464$ $\mathbf{0.524}$
-0.3 $0.652$ $0.708$ $1.056$ $0.528$ $0.564$
-0.5 $0.620$ $\mathbf{0.693}$ $1.227$ $0.475$ $0.561$
-0.7 $0.658$ - - - -
-SAGE 0.1 $0.735$ $0.726$ $\mathbf{0.739}$ $0.526$ $0.905$
-0.3 $0.671$ - - - -
-0.5 $0.695$ - - - -
-0.7 $0.645$ - - - -
-GIN 0.1 $0.644$ $0.726$ $0.746$ $0.494$ $0.856$
-0.3 $0.669$ - - - -
-0.5 $0.631$ - - - -
-0.7 $0.658$ - - - -
-Table 14: Results on Graph Regression (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
-Table 15: The Tables show the comparison between Extra Nodes and Cluster Nodes methods for node classification and node regression task.
-We show the average of the top 10 accuracies (higher the better) of the node classification task on Cora dataset and the average of the best 10 losses (lesser the better) of the node regression task on Chameleon dataset with Gs-train-to-Gs-infer experimental setup.
-They demonstrate the superior performance of the Cluster Nodes method over Extra Nodes.
-(a) Node Classification on Cora.
-r Extra Nodes Cluster Nodes
-$0.1$ $\mathbf{0.798\pm 0.008}$ $0.790\pm 0.004$
-$0.3$ $0.798\pm 0.008$ $\mathbf{0.800\pm 0.003}$
-$0.5$ $0.807\pm 0.004$ $\mathbf{0.829\pm 0.002}$
-$0.7$ $0.797\pm 0.007$ $\mathbf{0.813\pm 0.002}$
-(b) Node Regression on Chameleon.
-r Extra Nodes Cluster Nodes
-0.1 $\mathbf{0.504\pm 0.004}$ $0.506\pm 0.006$
-0.3 $0.579\pm 0.006$ $\mathbf{0.489\pm 0.003}$
-0.5 $0.585\pm 0.004$ $\mathbf{0.484\pm 0.004}$
-0.7 $\mathbf{0.504\pm 0.006}$ $0.539\pm 0.009$
-Table 16: Results comparing various coarsening methods on PROTEINS and ZINC (subset) datasets.
-The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
+variation_neighborhoods coarsening algorithm is used.
+Cora Chameleon
+Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
+variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
+algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
+kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
+heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
+variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
+variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
+Table 11: The table shows the FIT-GNN ablation study to compare various coarsening methods on Cora and Chameleon datasets, and the metrics for each dataset are accuracy (higher the better) and normalized MAE (lower the better), respectively.
 PROTEINS ZINC (subset)
@@
 variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
-Table 17: Results comparing various coarsening methods on Cora and Chameleon datasets and the metric for each dataset are accuracy (higher the better) and normalized MAE (lower the better) respectively.
-Cora Chameleon
-Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
-variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
-algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
-kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
-heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
-variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
-variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
+Table 12: The table shows the FIT-GNN ablation study to compare various coarsening methods on PROTEINS and ZINC (subset) datasets.
+The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
+Dataset Appending Nodes FIT-GNN Baseline
+$\mathbf{=0.1} r$$\mathbf{=0.3}$ $\mathbf{=0.5} r$$\mathbf{=0.7}$
+Chameleon Cluster Node $0.201$ $0.235$ $0.277$ $0.418$ $2.078$
+Extra Node $0.622$ $0.622$ $0.622$ $0.564$
+Crocodile Cluster Node $1.127$ $1.132$ $1.654$ $1.779$ $10.937$
+Extra Node $4.586$ $4.586$ $4.586$ $4.586$
+Squirrel Cluster Node $1.436$ $1.532$ $1.815$ $2.869$ $8.614$
+Extra Node $7.190$ $7.190$ $7.190$ $7.190$
+Cora Cluster Node $1.249$ $0.618$ $0.661$ $0.827$ $14.992$
+Extra Node $1.906$ $1.103$ $0.970$ $0.970$
+Citeseer Cluster Node $30.097$ $2.048$ $1.195$ $1.607$ $47.170$
+Extra Node $30.097$ $2.490$ $1.835$ $1.835$
+Pubmed Cluster Node $0.584$ $0.542$ $0.544$ $0.546$ $39.166$
+Extra Node $0.778$ $0.562$ $0.562$ $0.562$
+DBLP Cluster Node $5.785$ $2.803$ $1.761$ $1.875$ $112.514$
+Extra Node $28.271$ $6.008$ $2.308$ $2.207$
+Physics Coauthor Cluster Node $22.911$ $12.481$ $12.767$ $13.688$ $1115.079$
+Extra Node $41.066$ $19.936$ $19.936$ $19.936$
+OGBN-Products Cluster Node $28.399$ $2840.706$
+Table 13: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)

### v5

@@
 Moreover, it significantly reduces memory consumption by as much as $\mathbf{100\times}$ —thereby enabling scalable inference on large-scale graphs comprising millions of edges.
-We provide a theoretical characterization of the upper bound on the coarsening ratio under which our method’s time and space complexity remain superior to that of the standard approach.
+We provide a theoretical characterization of the upper bound on the expected size of subgraphs under which our method’s time and space complexity remain superior to that of the baselines.
 Notably, we observe that conventional methods often fail to perform inference on large graphs due to memory constraints, whereas our approach remains effective and scalable in such scenarios.
@@
 The aim of reducing computational costs has been approached through various methods.
-Research by 10.1145/3292500.3330925 and Zeng2020GraphSAINT: uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training.
-Other methods, including layer-wise sampling and mini-batch training, have been explored by chen2018stochastic; chen2018fastgcn; 10.1145/3394486.3403192; DBLP:journals/corr/abs-1911-07323.
-Additionally, DBLP:journals/corr/abs-2007-02817 introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while 10.1145/3447548.3467256 [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time.
-pmlr-v202-kumar23a; joly2024graph; NEURIPS2024_733209a1; 10.1145/3589335.3651920 also proposed different coarsening algorithms for scalable training approaches.
-jin2021graph [GCOND], wang2024fast [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph.
+Research by Chiang et al.
+(2019) and Zeng et al.
+(2020) uses a subgraph sampling technique, where a small subgraph is sampled at each training iteration for GNN training.
+Other methods, including layer-wise sampling and mini-batch training, have been explored by Chen et al.
+(2018a; b); Cong et al.
+(2020); Zou et al.
+(2019).
+Additionally, Fahrbach et al.
+(2020) introduced a coarsening algorithm based on Schur complements to create node embeddings in large graphs, while Huang et al.
+(2021) [SGGC] applied GNNs to coarsened graphs, reducing nodes and edges to lower training time.
+Kumar et al.
+(2023); Joly and Keriven (2024); Kataria et al.
+(2024); Dickens et al.
+(2024) also proposed different coarsening algorithms for scalable training approaches.
+Jin et al.
+(2021) [GCOND], Wang et al.
+(2024) [GC-SNTK] presented algorithms to generate a synthetic graph that mimics the original graph.
 However, their approach generates model-specific synthetic graphs that require training on the whole dataset, which defeats the purpose of reducing computational burden.
-gupta2025bonsai [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees.
+Gupta et al.
+(2025) [BONSAI] addressed this issue by looking into computational trees and training on just the relevant computational trees.
 However, these approaches do not decrease inference time since the inference phase takes the whole graph as input.
-DBLP:journals/corr/abs-2202-00075 shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach.
-In the graph-level task, Jin_2022 [DOSCOND] and 10.1145/3580305.3599398 [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK.
+Xue et al.
+(2023) shifted the training process to subgraph-level, using coarsening algorithms to split the graph into $k$-subgraphs and recommending a multi-device training approach.
+In the graph-level task, Jin et al.
+(2022) [DOSCOND] and Xu et al.
+(2023) [KIDD] tackled the graph classification problem following up on the approach of GCOND and GC-SNTK.
 However, the graph regression problem was not explored in this direction.
 3 Preliminaries
-3.1 Graph Notations
-Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix.
+Graph Notations: Given an undirected graph $G=(V,E,X,W)$, $V$ is the vertex set, $E$ is the edge set, $X\in\mathbb{R}^{n\times d}$ is the feature matrix, and $W$ is the edge weight matrix.
 Let $|V|=n$ be the number of nodes and $|E|=m$ be the number of edges.
-Let $A\in\{0,1\}^{n\times n}$ be the adjacency matrix of $G$, $d_{i}$ to be the degree of the node $v_{i}$.
+Let $A\in\mathbb{R}^{n\times n}$ be the adjacency matrix of $G$ where $A_{ij}$ is the edge weight between nodes $v_{i}$ and $v_{j}$, and $d_{i}$ is the degree of the node $v_{i}$.
 $D\in\mathbb{R}^{n\times n}$ is a diagonal degree matrix with $i$th diagonal element as $d_{i}$.
 We denote $\mathcal{N}_{j}(v_{i})$ as the $j$-hop neighbourhood of node $v_{i}\in V$.
-3.2 Graph Neural Network
-Graph Neural Network is a neural network designed to work for graph-structured data.
+Graph Neural Network: Graph Neural Network is a neural network designed to work for graph-structured data.
 The representation of each node in the graph is updated recursively by aggregation and transforming the node representations of its neighbors.
-kipf2017semisupervised introduced Graph Convolutional Network (GCN) as follows:
+Kipf and Welling (2017) introduced Graph Convolutional Network (GCN) as follows:
 $X^{(l+1)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(l)}\mathcal{W}^{(l)})$ (1)
@@
 $\mathcal{W}^{(l)}$ is the learnable parameter and $\sigma$ is a non-linear activation function.
-3.3 Graph Coarsening and Graph Partitioning
-loukas2018graphreductionspectralcut discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
-$X^{\prime}$ is the coarsened node representation.
+Graph Coarsening and Graph Partitioning: Loukas (2019) discussed multiple coarsening algorithms which create a coarsened graph $G^{\prime}=(V^{\prime},E^{\prime},X^{\prime},W^{\prime})$ from a given graph $G=(V,E,X,W)$.
 We refer the vertex set $V^{\prime}=\{v^{\prime}_{1},v^{\prime}_{2},\dots,v^{\prime}_{k}\}$ of $G^{\prime}$ as coarsened nodes.
-Given a coarsening ratio $r\in[0,1]$, we have $n\times r, kk\in\mathbb{Z}$.
+Given a coarsening ratio $r\in[0,1]$, we have $\lfloor{n\times r}\rfloor, kk\in\mathbb{Z}$.
 We can interpret it as creating $k$ disjoint partitions, $C_{1},C_{2},\dots,C_{k}$, from a graph of $n$ nodes.
 Mathematically, we create a partition matrix $P\in\{0,1\}^{n\times k}$, where $P_{ij}=1$ if and only if node $v_{i}\in C_{j}$.
-$A^{\prime}=P^{T}AP$ is the adjacency matrix of $G^{\prime}$.
-Similarly, the corresponding degree matrix $D^{\prime}=P^{T}DP$.
-SGGC (10.1145/3447548.3467256) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
+$X^{\prime}=P^{\top}X$ is the coarsened node representation.
+$A^{\prime}=P^{\top}AP$ is the adjacency matrix of $G^{\prime}$.
+Similarly, the corresponding degree matrix $D^{\prime}=P^{\top}DP$.
+SGGC (Huang et al., 2021) used normalized partition matrix $\mathcal{P}=PC^{-\frac{1}{2}}$, where $C$ is defined as a diagonal matrix with diagonal entries $C_{jj}=|C_{j}|,j=1,2,\dots,k$.
 Along with $G^{\prime}$, we create a set of disjoint subgraphs $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$ corresponding to the partitions $C_{1},C_{2},\dots,C_{k}$.
@@
 Each subgraph $G_{i}$ is the induced subgraph of $G$ formed by the nodes in $C_{i}$.
-$A_{i}$ and $D_{i}$ are the adjacency matrix and degree matrix of $G_{i}$ respectively.
+$A_{i}$ and $D_{i}$ are the adjacency and degree matrix of $G_{i}$ respectively.
 […]
@@
 In SGGC, a graph $G$ is reduced to $G^{\prime}$ using a partition matrix $P$ generated from the coarsening algorithm.
-Here $X^{\prime}=P^{T}X$.
-The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{T}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task.
+The labels for the coarsened graph are $Y^{\prime}=\arg\max(P^{\top}Y)$, where $Y$ is the label matrix of $G$ storing the one-hot encoding of the label of each node for the node classification task.
 Training is carried out on the coarsened graph $G^{\prime}$ using $Y^{\prime}$ as the label information.
@@
   • Extra Nodes: Adding the $1$-hop neighbouring nodes, namely Extra Nodes, in each subgraph $G_{i}$ denoted by $\mathcal{E}_{G_{i}}$.
-DBLP:journals/corr/abs-2202-00075 introduced this approach to reduce information loss after partitioning.
+Xue et al.
+(2023) introduced this approach to reduce information loss after partitioning.
 We add a unit weight edge if two nodes in $\mathcal{E}_{G_{i}}$ are connected in $G$.
@@
 Instead of adding the neighboring nodes, we will add Cluster Nodes, which represent the neighboring clusters, denoted by $\mathcal{C}_{G_{i}}$.
-liu2024scalable introduced this method to add nodes in the graph instead of adding to separate subgraphs.
+Liu et al.
+(2024) introduced this method to add nodes in the graph instead of adding to separate subgraphs.
 They also added cross-cluster edges among these Cluster Nodes and proposed a subgraph sampling-based method on the modified graph.
@@
 However, using Cluster Nodes, we only add one representative node for each neighboring cluster.
-Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is less than or equal to that with Extra Nodes.
+Hence, the time taken to train and infer these modified subgraphs with Cluster Nodes is at most that with Extra Nodes.
   • Lemma 4.1 shows that the Extra Nodes approach reduces the information loss due to partitioning for $1$ Layer of GNN.
@@
 2:Apply a coarsening algorithm on $G$, and output a normalized partition matrix $P$.
-3:Construct $\mathcal{G}_{s}=\{G_{1},G_{2},\dots,G_{k}\}$ using $P$;
+3:Construct $\mathcal{G}_{s}=\{G_{1}(V_{1},E_{1},X_{1}),G_{2}(V_{2},E_{2},X_{2}),\dots,G_{k}(V_{k},E_{k},X_{k})\}$ using $P$;
 4:if Append additional nodes then
-[…]
+5:  Append $\mathcal{E}_{G_{i}}$ or $\mathcal{C}_{G_{i}}$ to each $G_{i}$ $\triangleright$ Based on method of appending nodes
 6:end if
 7:Construct mask for each subgraph $Mask=\{mask_{1},mask_{2},\dots,mask_{k}\}$;
-8:Construct feature matrix $\{X_{1},X_{2},\dots,X_{k}\}$ for $\mathcal{G}_{s}$;
-9:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
+8:Construct label matrix $\{Y_{1},Y_{2},\dots,Y_{k}\}$ for $\mathcal{G}_{s}$ ;
 […]
-11:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
-12:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\quad\#$ Algorithm 4
-13:  $O.append(O_{i}[mask_{i}])$;
-14:  $Y.append(Y_{i}[mask_{i}])$;
-15:end for
+10:for each subgraph $G_{i}$ in $\mathcal{G}_{s}$ do
+11:  $O_{i}=M(L,A_{i},D_{i},X_{i})$; $\triangleright$ Algorithm 4
+12:  $O.append(O_{i}[mask_{i}])$;
+13:  $Y_{\textup{Collected}}.append(Y_{i}[mask_{i}])$;
+14:end for
 […]
-17:Train $M$ to minimize LossLoss;
+16:Train $M$ to minimize LossLoss;
 4.1 Node-level Task
@@
 On the other hand, our proposed Algorithm 1 uses the same partition matrix $P$ to construct and train on subgraphs $\mathcal{G}_{s}$ using original label information $Y$.
-For the node classification task, $\arg\max$ is used for creating labels for $G^{\prime}$ and CrossEntropy as a loss function.
+For the node classification task, $\arg\max$ is used for creating $Y^{\prime}$ for $G^{\prime}$ and CrossEntropy as a loss function.
 For the node regression task, we do not create $G^{\prime}$.
@@
 For a given graph $G\in\mathcal{D}$, we create $G^{\prime}$ and $\mathcal{G}_{s}=\{G_{1},G_{2},$ $\dots,G_{k}\}$.
-Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model to train on $G^{\prime}$ and Algorithm 2 to create a model to train on $\mathcal{G}_{s}$.
+Since $G^{\prime}$ is a single graph and $\mathcal{G}_{s}$ is a set of subgraphs, we use Algorithm 5 to create a model that trains on $G^{\prime}$ and Algorithm 2 to create a model that trains on $\mathcal{G}_{s}$.
 We use the CrossEntropy loss function for the graph classification task and the Mean Absolute Error (MAE) for the graph regression task.
@@
 4:  for $j=1$ to $L$ do
-[…]
+5:   $X_{i}^{(j)}=\sigma(\tilde{D_{i}}^{-\frac{1}{2}}\tilde{A_{i}}\tilde{D_{i}}^{-\frac{1}{2}}X_{i}^{(j-1)}\mathcal{W}^{(j-1)})$ $\triangleright$ Equation 1
 6:  end for
-[…]
+7:  $\tilde{X}.stack(X_{i}^{(L)})$ $\triangleright$ Stack along row axis
 8:end for
@@
 (a) Time Complexity
-Train Inference (Full Graph)
+Train Inference
 Classical $nd^{2}+n^{2}d$ $nd^{2}+n^{2}d$
@@
 FIT-GNN $\begin{matrix}k^{2}+kd+d^{2}\\ +\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]\end{matrix}$ $d^{2}+\max_{i=1}[\bar{n}_{i}^{2}+\bar{n}_{i}d]$
-In Equation 1, the matrix dimensions are as follows: $\tilde{A}$ has dimensions $(n\times n)$, $X^{(l)}$ is of dimensions $(n\times d)$, and $\mathcal{W}^{(l)}$ has dimensions $(d\times d)$.
-Time complexity for one layer of GNN computation is $\mathcal{O}(n^{2}d+nd^{2})$.
-If we take $L$ layers, then the total time is $\mathcal{O}(Ln^{2}d+Lnd^{2})$.
-The space complexity is $\mathcal{O}(n^{2}+Lnd+Ld^{2})$.
-When we compute on a sparse graph, the time complexity is $\mathcal{O}(m+Lnd^{2})$ and space complexity is $\mathcal{O}(m+Lnd+Ld^{2})$.
-The complexity is based on GCN.
-For other architectures like GAT veličković2018graphattentionnetworks, the complexity changes.
+[…]
 SGGC improved the time and space complexity (Table 1) for training the network by reducing the number of nodes from $n$ to $k$. However, the inference time and space complexity remain the same.
@@
 One is the classical model, where no coarsening or partitioning is done; the second is SGGC, where training is performed on a smaller graph $G^{\prime}$ and inference on $G$; the third is our approach, FIT-GNN, where both training and inference are performed on $\mathcal{G}_{s}$.
-The inference time complexity of our model is $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$, where $\bar{n}_{i}=n_{i}+\phi_{i}$ and $\phi_{i}$ is the number of additional nodes appended in each subgraph.
-Table 1 compares different approaches.
+[…]
 Lemma 4.2.
-The inference time complexity is at most $\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]\leq n^{2}d+nd^{2}$ if $\mathbb{E}[n_{i}+\phi_{i}]\leq\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\frac{d}{2}$.
+[…]
 Proof.
-$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=\sum_{i=1}^{k}[n_{i}^{2}d+\phi_{i}^{2}d+2n_{i}\phi_{i}d+n_{i}d^{2}+\phi_{i}d^{2}]$
-$=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
-Now, define $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$.
+We can expand our inference time complexity as follows:
+$\sum_{i=1}^{k}[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}]=nd^{2}+d\left[\sum^{k}_{i=1}n_{i}^{2}+\sum^{k}_{i=1}\phi_{i}^{2}+2\sum_{i=1}^{k}n_{i}\phi_{i}+d\sum_{i=1}^{k}\phi_{i}\right]$
+Now, define expected number of nodes in each subgraph $\mathbb{E}[n_{i}]:=\frac{1}{k}\sum_{i=1}^{k}n_{i}=\frac{n}{k}=\frac{1}{r}$ (since $k=nr$) and expected number of additional nodes appended in each subgraph $\mathbb{E}[\phi_{i}]:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}$.
 We also define the variances and covariance,
-$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2}$ $\qquad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},$
-$\text{Cov}(n_{i},\phi_{i})=\text{Cov}$ $:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
+$\text{Var}(n_{i}):=\frac{1}{k}\sum_{i=1}^{k}n_{i}^{2}-\mathbb{E}[n_{i}]^{2},\quad\text{Var}(\phi_{i}):=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}^{2}-\mathbb{E}[\phi_{i}]^{2},\quad\text{Cov}(n_{i},\phi_{i})=\text{Cov}:=\frac{1}{k}\sum_{i=1}^{k}\phi_{i}n_{i}-\mathbb{E}[\phi_{i}]\mathbb{E}[n_{i}]$
 We can also write,
@@
 If $\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$, then we can say the following:
-$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0$
-$\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$
+$\sqrt{\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})}-\left(\frac{d}{2}+\frac{1}{r}\right)\geq 0\Rightarrow\frac{d^{2}}{4}+\frac{d}{r}+\frac{n}{r}-\textup{Var}(n_{i}+\phi_{i})\geq\left(\frac{d}{2}+\frac{1}{r}\right)^{2}$
 $\Rightarrow\textup{Var}(n_{i}+\phi_{i})\leq\frac{n}{r}-\frac{1}{r^{2}}$
@@
   • Gc-train-to-Gs-train: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then use the learned weight as an initialization for subgraph-level training and final inference on $\mathcal{G}_{s}$.
-Our aim was to examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
+We examine whether pretraining on $G^{\prime}$ (which requires less time) followed by fine-tuning on $\mathcal{G}_{s}$ for fewer epochs can improve performance.
   • Gc-train-to-Gs-infer: Train the GNN model on $G^{\prime}$ as per Algorithm 3, then infer on $\mathcal{G}_{s}$ using the learned weights.
@@
 For the node regression task, we only perform Gs-train-to-Gs-infer because a coarsened graph is not created for node regression datasets.
-Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E. Source code: FIT-GNN.
+Dataset descriptions and details related to hyperparameters, device configuration, and key packages used for the experiments are in Section D and E.
+Source code: https://github.com/Roy-Shubhajit/FIT-GNN.
 Table 2: Datasets and corresponding splits used in our experiments.
 ‘c’ denotes the number of classes in node classification datasets.
-Public split “fixed” from 10.5555/3045390.3045396 is used for Cora, Citeseer, and PubMed.
+Public split “fixed” by Yang et al.
+(2016) is used for Cora, Citeseer, and PubMed.
 For the “random” split, we use 20 nodes per class for training, 30 per class for validation, and the rest for testing.
-For QM9, we follow DBLP:journals/corr/GilmerSRVD17 and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
+For QM9, we follow Gilmer et al.
+(2017) and predict one property from each broad category: dipole moment ($\mu$), HOMO-LUMO gap ($\epsilon_{\mathrm{HOMO}}-\epsilon_{\mathrm{LUMO}}$), zero-point vibration energy (ZPVE), and atomization energy ($U^{\mathrm{ATOM}}$) at $298.15\mathrm{K}$.
 Task Dataset (Reference) Train Val Test
-Node Regression Chameleon rozemberczki2019multiscale 30% 20% 50%
-Crocodile rozemberczki2019multiscale 30% 20% 50%
-Squirrel rozemberczki2019multiscale 30% 20% 50%
-Node Classification Cora McCallum2000 $20\times c$ $30\times c$ Rest
-Citeseer 10.1145/276675.276685 $20\times c$ $30\times c$ Rest
-Pubmed Sen_Namata_Bilgic_Getoor_Galligher_Eliassi-Rad_2008 $20\times c$ $30\times c$ Rest
-Coauthor Physics DBLP:journals/corr/abs-1811-05868 $20\times c$ $30\times c$ Rest
-DBLP 10.1145/1401890.1402008 $20\times c$ $30\times c$ Rest
-OGBN-Products 10.5555/3495724.3497579 $20\times c$ $30\times c$ Rest
-Graph Regression QM9 DBLP:journals/corr/WuRFGGPLP17 50% 25% 25%
-ZINC (Subset) DBLP:journals/corr/Gomez-Bombarelli16 50% 25% 25%
-Graph Classification PROTEINS Morris+2020 50% 25% 25%
-AIDS Morris+2020 50% 25% 25%
+Node Regression Chameleon (Rozemberczki et al., 2021) 30% 20% 50%
+Crocodile (Rozemberczki et al., 2021) 30% 20% 50%
+Squirrel (Rozemberczki et al., 2021) 30% 20% 50%
+Node Classification Cora (McCallum et al., 2000) $20\times c$ $30\times c$ Rest
+Citeseer (Giles et al., 1998) $20\times c$ $30\times c$ Rest
+Pubmed (Sen et al., 2008) $20\times c$ $30\times c$ Rest
+Coauthor Physics (Shchur et al., 2018) $20\times c$ $30\times c$ Rest
+DBLP (Tang et al., 2008) $20\times c$ $30\times c$ Rest
+OGBN-Products (Hu et al., 2020) $20\times c$ $30\times c$ Rest
+Graph Regression QM9 (Wu et al., 2018) 50% 25% 25%
+ZINC (Subset) (Gómez-Bombarelli et al., 2018) 50% 25% 25%
+Graph Classification PROTEINS (Morris et al., 2020) 50% 25% 25%
+AIDS (Morris et al., 2020) 50% 25% 25%
 6 Result and Analysis
 6.1 Model Performance
+Table 3: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$. variation_neighborhoods coarsening algorithm is used. OOM means Out Of Memory.
+Full (Luo et al., 2024) $0.823\pm 0.001$
+SGGC OOM
+GCOND OOM
+BONSAI OOM
+FIT-GNN $\mathbf{0.894\pm 0.000}$
+Table 4: Results for node classification tasks with accuracy as the metric (higher is better).
+We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
+variation_neighborhoods coarsening algorithm is used.
+Methods Models Reduction Ratio (r) Dataset
+Cora Citeseer Pubmed DBLP Physics
+Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $\mathbf{0.933\pm 0.000}$
+GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
+SGGC GCN 0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
+0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
+GAT 0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
+0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
+GCOND GCN 0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
+0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
+GAT 0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
+0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
+BONSAI GCN 0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
+0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
+GAT 0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
+0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
+FIT-GNN GCN 0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $0.932\pm 0.000$
+0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
+GAT 0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
+0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
+Figure 3: The plot shows the ablation study conducted on the Cora dataset to determine which experimental setup performs better than the others.
+The plot also compares different methods of appending nodes to the subgraphs and how performance changes with varying coarsening ratios.
+variation_neighborhoods coarsening algorithm is used.
 Table 5: This table shows the normalized MAE loss (lower is better) for node regression task on different datasets with Gs-train-to-Gs-infer experiment setup and Cluster Nodes method.
@@
 Methods Models Reduction Ratio (r) Dataset
-Chameleon Crocodile Squirel
+Chameleon Crocodile Squirrel
 Full GCN 1.0 $0.852\pm 0.001$ $0.854\pm 0.000$ $0.802\pm 0.000$
@@
 0.7 $0.792\pm 0.089$ $0.779\pm 0.102$ $0.791\pm 0.009$
-Table 4: Results for node classification tasks with accuracy as the metric (higher is better).
-We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
-variation_neighborhoods coarsening algorithm is used.
-Methods Models
-Reduction
-Ratio (r)
-Dataset
-Cora Citeseer Pubmed DBLP Physics
-Full GCN 1.0 $0.821\pm 0.002$ $0.706\pm 0.002$ $0.768\pm 0.020$ $0.724\pm 0.002$ $\mathbf{0.933\pm 0.000}$
-GAT 1.0 $0.809\pm 0.004$ $0.700\pm 0.003$ $0.740\pm 0.004$ $0.713\pm 0.004$ $0.909\pm 0.008$
-SGGC GCN 0.3 $0.808\pm 0.003$ $0.700\pm 0.001$ $0.773\pm 0.002$ $0.737\pm 0.011$ $0.928\pm 0.003$
-0.5 $0.808\pm 0.001$ $0.716\pm 0.002$ $\mathbf{0.793\pm 0.002}$ $0.733\pm 0.008$ $0.931\pm 0.004$
-GAT 0.3 $0.686\pm 0.010$ $0.635\pm 0.021$ $0.773\pm 0.002$ $0.626\pm 0.028$ $0.870\pm 0.016$
-0.5 $0.641\pm 0.026$ $0.635\pm 0.024$ $0.683\pm 0.024$ $0.604\pm 0.047$ $0.842\pm 0.018$
-GCOND GCN 0.3 $0.806\pm 0.003$ $0.722\pm 0.001$ $0.779\pm 0.002$ $0.748\pm 0.001$ $0.876\pm 0.011$
-0.5 $0.808\pm 0.004$ $\mathbf{0.730\pm 0.007}$ $0.773\pm 0.004$ $0.755\pm 0.002$ $0.804\pm 0.005$
-GAT 0.3 $0.549\pm 0.080$ $0.628\pm 0.095$ $0.370\pm 0.079$ $0.481\pm 0.034$ OOM
-0.5 $0.712\pm 0.059$ $0.501\pm 0.085$ $0.409\pm 0.044$ $0.464\pm 0.030$ OOM
-BONSAI GCN 0.3 $0.722\pm 0.011$ $0.585\pm 0.003$ $0.538\pm 0.103$ $0.683\pm 0.039$ $0.798\pm 0.007$
-0.5 $0.701\pm 0.012$ $0.666\pm 0.001$ $0.670\pm 0.084$ $0.702\pm 0.049$ $0.832\pm 0.016$
-GAT 0.3 $0.680\pm 0.014$ $0.566\pm 0.003$ $0.540\pm 0.096$ $0.665\pm 0.027$ $0.791\pm 0.010$
-0.5 $0.683\pm 0.006$ $0.579\pm 0.004$ $0.523\pm 0.133$ $0.681\pm 0.025$ $0.812\pm 0.016$
-FIT-GNN GCN 0.3 $0.800\pm 0.003$ $0.679\pm 0.003$ $0.754\pm 0.001$ $\mathbf{0.786\pm 0.001}$ $0.932\pm 0.000$
-0.5 $\mathbf{0.829\pm 0.002}$ $0.668\pm 0.003$ $0.761\pm 0.004$ $0.700\pm 0.001$ $0.926\pm 0.000$
-GAT 0.3 $0.761\pm 0.004$ $0.655\pm 0.005$ $0.754\pm 0.002$ $0.771\pm 0.002$ $0.885\pm 0.004$
-0.5 $0.792\pm 0.003$ $0.669\pm 0.004$ $0.760\pm 0.003$ $0.720\pm 0.002$ $0.885\pm 0.003$
-Node Regression: Table 3 shows the node regression error (mean and standard deviation of the lowest $10$ MAE).
-As the coarsening ratio increases, more subgraphs are created, and the size of each subgraph also reduces, and important neighboring node information gets lost.
-Hence, an increase in the coarsening ratio decreases performance.
-Table 5: Results on OGBN-Products. FIT-GNN model uses a coarsening ratio $0.5$. variation_neighborhoods coarsening algorithm is used. OOM means Out Of Memory.
-Full $0.873\pm 0.000$
-SGGC OOM
-GCOND OOM
-BONSAI OOM
-FIT-GNN $\mathbf{0.894\pm 0.000}$
 Node Classification: Table 4 shows the results for node classification accuracy (mean and standard deviation of the top $10$ accuracies) with coarsening ratios $0.3$ and $0.5$ using the Cluster Nodes approach.
@@
 From the Figure, we observe that Gc-train-to-Gs-train offers the best result among the three experimental setups for the node classification task.
-For the ’None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance.
-We also observe in the ’None’ method that, with an increase in the coarsening ratio, the performance drops.
+For the ‘None’ method of appending nodes, where no additional nodes are added to each subgraph, we notice that it performs poorly compared to other methods, indicating that the Extra Node and Cluster Node methods indeed provide additional information for better empirical performance.
+We also observe in the ‘None’ method that, with an increase in the coarsening ratio, the performance drops.
 This is because, as the coarsening ratio increases, the size of each subgraph gets smaller, hence more edges are cut due to partitioning, leading to more information loss.
 For the Gc-train-to-Gs-infer method, we observe that the performance drops as the coarsening ratio increases.
-This can be argued because of the change in the distribution of the coarsening graph and the set of subgraphs.
+This can be argued because of the change in the distribution of the coarsened graph and the set of subgraphs.
 When training on $G_{s}$, the distribution in the train and test sets is similar, hence attaining better performance.
@@
 Table 3 shows results on the OGBN-Products dataset.
-Result on training on the full dataset from the OGBN-Products leaderboard [Link].
-For training on the full dataset, $8\%$ for training, next top $2\%$ for validation, and the rest for testing.
-For our approach, we use $0.038\%$ training, $0.057\%$ validation, and the rest for testing.
+We use the result by Luo et al.
+(2024) for the full graph setting, which used the OGBN’s standard split, while the split used in our work is detailed in Table 2.
+Node Regression: Table 5 presents the node regression error (mean and standard deviation of the lowest 10 MAE).
+Interestingly, our results demonstrate that utilizing localized subgraphs for inference (FIT-GNN) significantly improves performance compared to full-graph inference.
+To investigate this counterintuitive performance leap, we conducted a rigorous ablation study (Appendix G).
+Our analysis indicates that this improvement is driven by the structural properties of the input data during inference rather than the training regime itself.
+First, partitioning the graph creates localized contexts that are statistically more homogeneous; we observed that the label standard deviation within individual subgraphs is drastically lower than the global variation, presenting a simpler optimization landscape for the model.
+Second, while coarsening (e.g., at a ratio of $r=0.5$) causes a vast majority of nodes to lose a significant portion of their distant 2nd-hop neighborhood, this structural loss actually acts as implicit adversarial pruning.
+In these specific heterophilic graphs, long-range information often introduces noise or adversarial signals.
+By filtering out this noise through coarsening, the model is able to fully exploit the low-variance local structures, ultimately leading to the observed reduction in regression error.
 Table 6: Results on Graph Regression with MAE as metric (lower is better); We use Extra node and Gs-train-to-Gs-infer as experimental setup. variation_neighborhoods coarsening algorithm is used.
 Methods Models Reduction Ratio (r) Dataset
-ZINC (Subset) QM9 (prop 0) QM9 (prop 4) QM9 (prop 6) QM9 (prop 13)
+ZINC (Subset) QM9 ($\mu$) QM9 ($\Delta\epsilon$) QM9 (ZPVE) QM9 ($U^{\textup{ATOM}}$)
 Full GCN 1.0 $0.743$ $0.855$ $1.012$ $1.102$ $1.073$
@@
 It is also observed that a lower coarsening ratio yields better loss, which implies that the model performs better on molecular graphs when the subgraph size is large.
-It implies that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction.
-Similarly, when inferred on a coarsened graph for each graph in the dataset, the performance degrades.
-This is because, when a graph is reduced to a smaller graph, the positional information is lost, which is crucial since in graphs representing different molecules, the position of a certain component in a molecule results in different properties of the molecule.
+We observe that finer subgraphs lose out on global information about the molecule, which is necessary for the prediction.
+Similarly, the performance degrades in the Gc-train-to-Gc-infer setup, where we infer on the coarsened graphs.
+This is because, when a graph is reduced to a smaller graph, the individual node information is lost, which is crucial since, in graphs representing different molecules, components formed by different atoms result in different properties of the molecule.
+Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$.
+For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data.
+DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution.
+Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class.
+However, the graph size is kept similar to the graph size in the training set.
+These approaches also don’t cover the test set.
+Compared to these baselines, our approach transforms each graph into a set of subgraphs and a coarsened graph based on a reduction ratio.
+As a result, each graph in the training set and test set gets reduced.
+From our results, we see that FIT-GNN outperforms all the baselines.
+Table 14 and 15 present the ablation study on various coarsening algorithms for all the aforementioned tasks.
+From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others.
 Table 7: Results on Graph Classification with accuracy as metric (higher is better).
@@
 50 $0.634$ $0.674$
-Methods Models
-Reduction
-Ratio (r)
-Dataset
+Methods Models Reduction Ratio (r) Dataset
 AIDS PROTEINS
@@
 0.7 $0.819$ $0.739$
-Graph Classification: Table 7 shows graph classification results on AIDS and PROTEINS datasets, where training and inference are done on $G^{\prime}$.
-For baselines like DOSCOND and KIDD, we use ‘Graph per class’, whereas we use ‘Reduction Ratio (r)’ as a metric to denote how these algorithms reduce the size of training data.
-DOSCOND and KIDD propose to generate synthetic graphs that mimic the training data distribution.
-Therefore, to cover all the graphs in the dataset appropriately, this approach generates {1, 10, 50} graphs per class.
-However, the graph size is kept similar to the graph size in the training set.
-These approaches also don’t cover the test set.
-Compared to these baselines, our approach reduces each graph into a set of subgraphs and a coarsened graph based on a reduction ratio.
-As a result, each graph in the training set and test set gets reduced.
-From our results, we see that FIT-GNN outperforms all the baselines.
-Table 11 and 12 present the ablation study on various coarsening algorithms for all the aforementioned tasks.
-From the results, we observe that variation_neighborhoods coarsening algorithm yields consistently better results compared to others.
-Therefore, we report results on all datasets using variation_neighborhoods coarsening algorithm.
 6.2 Inference Time and Memory
@@
 We show single-node prediction time using two different coarsening ratios for Node-Level tasks.
-Here, we used Cluster Nodes for the node-level task.
+Here, we used Cluster Nodes.
 We used Gc-train-to-Gc-infer experimental setup for the graph-level tasks.
+(a) Node-Level Tasks
 FIT-GNN
@@
 OGBN-Products $0.1762$ $0.0017$ $\mathbf{0.0016}$
-(a) Node-Level Tasks
+(b) Graph-Level Tasks
 FIT-GNN
-Dastaset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
+Dataset Baselines $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
 ZINC (subset) $\mathbf{0.00184}$ $\mathbf{0.00184}$ $0.00190$
@@
 PROTEINS $0.00165$ $\mathbf{0.00160}$ $0.00163$
-(b) Graph-Level Tasks
 […]
@@
 In contrast, the FIT-GNN model only requires the relevant subgraph, resulting in faster predictions.
-For larger datasets like the subset of OGBN-Products, baseline inference is not feasible with our computational capacity.
-Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges.
-We see up to $\mathbf{100\times}$.
+For larger datasets like OGBN-Products, baseline inference is not feasible with our computational capacity.
+Therefore, to show a comparison, we take a subset of OGBN-Products with $165000$ nodes and $4340428$ edges with respect to which we see up to $\mathbf{100\times}$ speedup in inference time.
 Section C elaborates more on the inference time for datasets with a larger number of nodes.
 Traditional methods fail to infer for very large datasets due to memory constraints.
-However, our approach can give an inference on the whole graph.
+However, our approach enables inference on the whole graph.
 Figure 4 and Table 13 show the detailed comparison of memory consumption for different datasets for both Extra Nodes and Cluster Nodes methods, along with the baseline memory, which highlights that the FIT-GNN model uses up to $\mathbf{100\times}$ less memory than the Baseline.
 Figure 6 shows the time taken to coarsen for different coarsening ratios for the Cora dataset.
-As the coarsening ratio increases, the time also increases, since the number of subgraphs increases with the coarsening ratio.
+As the coarsening ratio increases, the time also increases, since the number of subgraphs increases.
 Table 8(b) compares the inference time of graph classification and graph regression tasks.
@@
 Empirically, we have shown that our method is comparable to the uncoarsened Baseline while being orders of magnitude faster in terms of inference time and consuming a fraction of the memory.
-A few possible future directions involve studying directed and weighted graphs, focusing on the theoretical connections between Extra Nodes and Cluster Nodes, and tackling real-world data-intensive challenges, for example, weather forecasting.
+Future directions include extending the FIT-GNN methodology to directed, weighted, heterogeneous, and temporal graphs, alongside tackling data-intensive real-world challenges like weather forecasting.
+Theoretically, we aim to quantify the information loss for the Cluster Node method, investigate its connection to Extra Nodes, and explore novel strategies beyond both to further mitigate graph partitioning loss.
+8 Acknowledgment
+Anirban Dasgupta acknowledges the support by SERB MATRICS and SERB CRG grants and the support from N Rama Rao Chair Professorship.
 Appendix
@@
 To understand the information loss when taking $2$ Layers of GNN, we divide $G_{i}$ into $3$ sets of nodes.
-  • Nodes with 1-hop and 2-hop neighbours in $G_{i}$. Let us call this set $S_{1}$
-  • Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{2}$
-  • Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$. Let us call this set $S_{3}$.
+  • $S_{1}$: Nodes with 1-hop and 2-hop neighbours in $G_{i}$.
+  • $S_{2}$: Nodes with 1-hop neighbours in $G_{i}$ but $\exists v$ in 2-hop neighbourhood that is not in $G_{i}$.
+  • $S_{3}$: Node where $\exists v$ in 1-hop and 2-hop neighbourhood that is not in $G_{i}$.
 Let $\mathcal{I}^{2}_{i}$ be the number of nodes whose information doesn’t get passed on after $2$ layer of GNN for $G_{i}$.
@@
 A.2 Cluster Nodes
-Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{T}X$.
+Given a partition matrix $P$, the features of the coarsened node are $X^{\prime}=P^{\top}X$.
 Given a normalized partition matrix, the features of a node $v^{\prime}_{i}\in G^{\prime}$ is the degree-weighted average of the features of nodes in $C_{i}$.
@@
 Appendix B More Algorithms
-Algorithm 3 Training GNN (Train on $G^{\prime}$) (10.1145/3447548.3467256)
+Algorithm 3 Training GNN (Train on $G^{\prime}$) (Huang et al., 2021)
 1:$G=(V,E,X)$; Labels $Y$; Node-Model $M$; Loss $\ell$; Number of Layers $L$
@@
 3:Construct Coarsened graph $G^{\prime}$ using $P$;
-4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{T}X$;
-5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{T}Y)$ ;
+4:Construct feature matrix for $G^{\prime}$ by $X^{\prime}=P^{\top}X$;
+5:Construct labels for $G^{\prime}$ by $Y^{\prime}=\arg\max(P^{\top}Y)$ ;
 […]
@@
 2:for $i=1$ to $L$ do
-3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\quad\#$ Equation 1
+3:  $X^{(i)}=\sigma(\tilde{D}^{-\frac{1}{2}}\tilde{A}\tilde{D}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})$ $\triangleright$ Equation 1
 4:end for
@@
 2:for $i=1$ to $L$ do
-[…]
+3:  $X^{(i)}$$\sigma(\tilde{D^{\prime}}^{-\frac{1}{2}}\tilde{A^{\prime}}\tilde{D^{\prime}}^{-\frac{1}{2}}X^{(i-1)}\mathcal{W}^{(i-1)})\triangleright$ Equation 1
 4:end for
@@
 Appendix C More on Time and Space complexity
+C.1 Comparison with Baseline
 In Lemma 4.2, we provide the conditions that need to be satisfied for better space and time complexities.
@@
     $\underset{\textup{\tiny Baseline Complexity}}{\underbrace{n^{2}d+nd^{2}}}\geq\underset{\textup{ \tiny FIT-GNN Complexity}}{\underbrace{\sum_{i=1}^{k}\bigl[(n_{i}+\phi_{i})^{2}d+(n_{i}+\phi_{i})d^{2}\bigr]}}$ (5)
-Figure 4: The figure shows the feasibility of our inference methods for different node-level datasets against coarsening ratios.
-Two important inference setups for each dataset: single-node inference with the computational cost $\mathcal{O}\left(\max_{i}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})+n\right)$, and full-graph inference with computational cost $\mathcal{O}\left(\sum_{i=1}^{k}(\bar{n_{i}}^{2}d+\bar{n_{i}}d^{2})\right)$, are compared against the baseline (classical GNN) computational cost of $\mathcal{O}\left(n^{2}d+nd^{2}\right)$.
-The coarsening method used is variation_neighborhoods.
-Both axes are in logarithmic scale (base 10).
+[…]
 Figure 5 demonstrates the feasibility of our FIT-GNN method for different coarsening ratios with regards to full-graph inference time complexity.
@@
   3. 3.
-    We observe lower inference time complexity of Cluster_node method as compared to Extra_node method under both inference setups.
+    We observe lower inference time complexity of Cluster Node method as compared to Extra Node method under both inference setups.
 This is expected as $|C_{G_{i}}|\leq|\mathcal{E}_{{G_{i}}}|,\ \forall i\in\{1,2,\dots,k\}$.
+C.2 Preprocessing Time Complexity
+A comparative analysis of our preprocessing (coarsening) overhead against other state-of-the-art scaling methods has been conducted.
+As shown in the Table 9 below, our preprocessing time complexity is comparable to, and often faster than, existing SOTA methods.
+Thus, even when end-to-end latency is considered, FIT-GNN maintains its efficiency advantage.
+Table 9: Asymptotic time complexities for different methods.
+NN is the number of nodes in the original graph, $M$ is the number of edges and $C$ is the number of classes; rest of the symbols follow notation defined in Section 3 and 4.
+Blue and Red indicate CPU and GPU respectively - the hardware used for computation.
+Full Graph setup indicates testing on all the test nodes in a graph while Single Node setup represents testing for a single test node.
+Method Preprocessing Training Inference (Full Graph) Inference (Single Node) Model-Agnostic
+SGGC $M+N$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$
+GCOND $\begin{matrix}C(N^{2}+k^{2})d\\ +C(N+k)d^{2}\end{matrix}$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$ $\times$
+BONSAI $M+N$ $k^{2}d+kd^{2}$ $N^{2}d+Nd^{2}$ $N^{2}d+Nd^{2}$
+FIT-GNN $M+N$ $\begin{matrix}kd^{2}+k^{2}d\\ +\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]\end{matrix}$ $\sum_{i=1}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$ $\max_{i}^{k}[\bar{n}_{i}^{2}d+\bar{n}_{i}d^{2}]$
+We further extend our complexity analysis to the practical scenario, where a new test node $v$ is introduced to the graph $G$.
+This setting is critical for real-world applications where graphs evolve dynamically.
+We compare three distinct inference strategies for this new node:
+  1. 1.
+    Full Graph: Construct $G_{new}$ by adding $v$ to the original full graph $G$ and infer on $G_{new}$ using the pre-trained network.
+  2. 2.
+    2nd-hop Neighborhood: Sample only the 2nd-hop neighborhood of $v$ from $G$ and infer using the pre-trained network (assuming a 2-layer architecture).
+  3. 3.
+    FIT-GNN Subgraph: Assign $v$ to a relevant subgraph $G_{i}$ (e.g., the one containing the majority of its 1st-hop neighborhood) and append the necessary Extra/Cluster nodes.
+Inference is then performed strictly within the modified subgraph $G_{i}$.
+We present the time complexity comparison for these methods in the Table 10 below.
+The most significant advantage of the FIT-GNN approach is that inference complexity depends solely on the subgraph size $\bar{n}_{i}$ ($\bar{n}_{i}\ll n$).
+In contrast, the Full Graph approach remains tied to the global parameter $n$ due to the necessity of accessing the full adjacency matrix.
+Table 10: Time complexity of different inference strategies when a new node $v$ is added to graph $G$
+Inference Strategy Preprocessing Complexity Inference Complexity Notes
+[…]
+[…]
+[…]
 Appendix D Dataset Description
+(a) Summary of datasets used for Graph classification
 Dataset
@@
 AIDS $2000$ 77 $16$ $38$ 22
-(a) Summary of datasets used for Graph classification Dataset
+(b) Summary of datasets used for Graph regression
 Dataset
@@
 ZINC (subset) $10000$ $11$ $25$ 11 11
-(b) Summary of datasets used for Graph regression Dataset
+(c) Summary of datasets used for Node classification
 Dataset Nodes Edges Features Classes
@@
 OGBN-Products $2449029$ $61859140$ $100$ $47$
-(c) Summary of datasets used for Node classification Dataset
+(d) Summary of datasets used for Node regression
 Dataset Nodes Edges Features
@@
 Crocodile $11631$ $170845$ $128$ 11
-(d) Summary of datasets used for Node regression Dataset
 Appendix E Experiment Details (Parameters and Device Configuration)
 For node classification and node regression tasks, we use Adam Optimizer with a learning rate of $0.01$ and $L_{2}$ regularization with $0.0005$ weight.
-We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph level tasks.
-For both coarsened graph (Gc-train) and set of subgraph based (Gs-train) model, we set epochs to 300 and the number of layers of GCN to 2 for training.
+We use Adam Optimizer with a learning rate of $0.0001$ and $L_{2}$ regularization with $0.0005$ weight for graph-level tasks.
+We set epochs to 20 and the number of layers of GCN to 2 for training.
 We set the hidden dimensions for each layer of GCN to $512$.
-The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to to train our models.
+The device configurations are Intel(R) Xeon(R) Gold 5120 CPU @ 2.202.20GHz, 256256GB RAM, NVIDIA A100 4040GB GPU. We use Pytorch Geometric to train our models.
 Appendix F Results
-Figure 5 presents the ablation study conducted to compare the performance of FIT-GNN for different experimental setup using different methods for appending nodes in the subgraph.
-The experiments are conducted on the Cora dataset for different coarsening ratios.
-Figure 5: The plot shows the ablation study conducted on the Cora dataset to determine which experimental setup performs better than the others.
-The plot also compares different methods of appending nodes to the subgraphs and how performance changes with varying coarsening ratios.
+Table 12 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$.
+We did not compare our experiments with methods such as Cluster-GCN(Chiang et al., 2019) because the method focuses on reducing training time by sampling subgraphs for several iterations.
+During the inference phase, however, it still performs inference on the entire graph.
+Consequently, its inference performance would be at most comparable to that of a standard GCN.
+Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios.
+It is observed that the None method takes the least amount of time, which is intuitive since the subgraphs are not appended with additional nodes.
+For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended.
+For Cluster Node, we observe the maximum time taken since we are not only looking at the neighboring nodes that are lost due to partition, but also are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
+Table 12: Results for node classification tasks with accuracy as the metric (higher is better).
+We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
 variation_neighborhoods coarsening algorithm is used.
-Figure 6 shows the time comparison for creating the subgraphs for Cora Dataset for different coarsening ratios.
-It is observed that the None method takes the least amount of time, which is obvious since the subgraphs are not appended with additional nodes.
-For the Extra Node Method, the time increases since the subgraph size increases as new nodes are appended.
-For Cluster Node, we observe the maximum time taken since not only are we looking at the neighboring nodes that are lost due to partition, but also we are keeping a track of the cluster to which these neighboring nodes belong, corresponding to which, we add the cross-cluster edges.
-Figure 6: The plot shows how the coarsening time varies for the Cora dataset for different coarsening ratios using different methods of appending nodes to the subgraphs.
-variation_neighborhoods coarsening algorithm is used.
-Table 10 presents the results on the node classification task with all the coarsening ratios $\{0.1,0.3,0.5,0.7\}$.
-To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 11 and 12.
-Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes.
-Note that this memory consumption is to store the graph and the weight matrix in the memory only.
 Methods Models Reduction Ratio (r) Dataset
@@
 0.7 $0.797\pm 0.003$ $0.662\pm 0.003$ $0.759\pm 0.002$ $0.739\pm 0.002$ $0.903\pm 0.003$
-Table 10: Results for node classification tasks with accuracy as the metric (higher is better).
-We use the Cluster Nodes method to append additional nodes to subgraphs and Gs-train-to-Gs-infer as experimental setup.
+Figure 6: The plot shows how the coarsening time varies for the Cora dataset for different coarsening ratios using different methods of appending nodes to the subgraphs.
 variation_neighborhoods coarsening algorithm is used.
-Cora Chameleon
-Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
-variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
-algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
-kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
-heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
-variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
-variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
-Table 11: The table shows the FIT-GNN ablation study to compare various coarsening methods on Cora and Chameleon datasets, and the metrics for each dataset are accuracy (higher the better) and normalized MAE (lower the better), respectively.
-PROTEINS ZINC (subset)
-Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
-variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.573}$ $\mathbf{0.620}$
-algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
-kron $0.522$ $0.652$ $0.663$ $1.201$
-heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
-variation_edges $0.652$ $0.565$ $1.548$ $1.883$
-variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
-Table 12: The table shows the FIT-GNN ablation study to compare various coarsening methods on PROTEINS and ZINC (subset) datasets.
-The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
+Table 13 shows the maximum GPU memory consumption during inference for different datasets with different coarsening ratios and methods of appending nodes.
+Note that this memory consumption is to store the graph and the weight matrix in the memory only.
+Table 13: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)
 Dataset Appending Nodes FIT-GNN Baseline
@@
 OGBN-Products Cluster Node $28.399$ $2840.706$
-Table 13: Summary of maximum GPU memory consumption during inference of datasets used for node-level tasks. All units are in MegaBytes (MB)
+To validate the use of “variation_neighborhoods” as a coarsening algorithm for all the results presented, we show an ablation study for different coarsening algorithms in Table 14 and 15.
+Table 14: The table shows the FIT-GNN ablation study to compare various coarsening methods on Cora and Chameleon datasets, and the metrics for each dataset are accuracy (higher the better) and normalized MAE (lower the better), respectively.
+Cora Chameleon
+Coarsening Method $\mathbf{r=0.1}$ $\mathbf{r=0.3}$ $\mathbf{r=0.1}$ $\mathbf{r=0.3}$
+variation_neighborhoods $0.790\pm 0.004$ $0.800\pm 0.003$ $\mathbf{0.496\pm 0.005}$ $\mathbf{0.489\pm 0.003}$
+algebraic_JC $\mathbf{0.827\pm 0.006}$ $\mathbf{0.804\pm 0.003}$ $0.544\pm 0.005$ $0.504\pm 0.003$
+kron $0.758\pm 0.004$ $0.800\pm 0.002$ $0.571\pm 0.004$ $0.580\pm 0.013$
+heavy_edge $0.736\pm 0.012$ $0.773\pm 0.006$ $0.572\pm 0.002$ $0.531\pm 0.002$
+variation_edges $0.484\pm 0.006$ $0.471\pm 0.012$ $0.513\pm 0.003$ $0.542\pm 0.005$
+variation_cliques $0.751\pm 0.012$ $0.801\pm 0.009$ $0.674\pm 0.009$ $0.679\pm 0.008$
+Table 15: The table shows the FIT-GNN ablation study to compare various coarsening methods on PROTEINS and ZINC (subset) datasets.
+The metric for the PROTEINS dataset is accuracy (higher is better), and for the ZINC (subset) is normalized MAE (lower is better)
+PROTEINS ZINC (subset)
+Coarsening Method $\mathbf{r=0.3}$ $\mathbf{r=0.5}$ $\mathbf{r=0.3}$ $\mathbf{r=0.5}$
+variation_neighborhoods $0.652$ $\mathbf{0.739}$ $\mathbf{0.573}$ $\mathbf{0.620}$
+algebraic_JC $\mathbf{0.826}$ $\mathbf{0.739}$ $0.823$ $0.810$
+kron $0.522$ $0.652$ $0.663$ $1.201$
+heavy_edge $0.565$ $0.609$ $2.114$ $1.227$
+variation_edges $0.652$ $0.565$ $1.548$ $1.883$
+variation_cliques $0.696$ $0.652$ $0.689$ $0.742$
+Appendix G Detailed Ablation Study on Node Regression Performance
+To understand the counterintuitive performance gains observed in node regression tasks when using FIT-GNN, we conducted a comprehensive ablation study.
+Specifically, we isolated the source of the performance gap to determine why restricting inference to localized subgraphs outperforms full-graph inference.
+G.1 Impact of Inference Input vs. Training Regime
+Our first objective was to determine whether the performance gain stemmed from the training methodology itself or the structure of the input data (subgraphs) during inference.
+We evaluated two distinct setups on the Crocodile dataset using a GCN convolutional layer:
+  • Setup A: Train on subgraphs $\rightarrow$ Infer on the full graph.
+  • Setup B: Train on the full graph $\rightarrow$ Infer on the full graph.
+Table 16: Comparison of training and inference setups to isolate the source of performance gains.
+Train Setup Inference Setup Mean Absolute Error (MAE)
+Full Graph Full Graph 0.852
+Subgraphs Full Graph 0.865
+Subgraphs (FIT-GNN) Subgraphs 0.364
+As shown in Table 16, the performance of Setup A and Setup B is nearly identical, indicating that training on subgraphs alone does not account for the drastic MAE reduction.
+The substantial performance leap in FIT-GNN only occurs when subgraphs are used as the inference input, confirming that the local structural input drives the improvement.
+G.2 Subgraph Optimization Landscape
+We further hypothesized that partitioning the graph into subgraphs presents a simpler optimization landscape for the model.
+To quantify this, we compared the variation of labels locally (within individual subgraphs) versus globally across the entire graph.
+We computed the Label Standard Deviation for node regression datasets and Label Entropy for node classification datasets.
+Table 17: Comparison of global label variation versus average local subgraph variation.
+Dataset Metric Used Global Variation Subgraph Variation (Avg)
+Cora Entropy 1.8311 0.1245
+Citeseer Entropy 1.7533 0.1572
+Chameleon Standard Deviation 2.1329 0.0689
+Squirrel Standard Deviation 1.7639 0.1284
+The results in Table 17 demonstrate that the variation of labels within individual subgraphs is drastically lower than the global variation.
+By coarsening the graph, we create local contexts that are statistically more homogeneous, allowing the GNN to specialize more effectively and navigate a smoother optimization landscape during inference.
+G.3 Structural Information Loss and Implicit Adversarial Pruning
+Finally, we investigated the role of structural information loss. Using a coarsening ratio of $r=0.5$, we empirically computed the fraction of the 2nd-hop neighborhood lost for each node.
+(a) Citeseer (Classification)
+(b) Cora (Classification)
+(c) Squirrel (Regression)
+(d) Chameleon (Regression)
+Figure 7: Histograms detailing the fraction of the 2nd-hop neighborhood lost for each node at a coarsening ratio of $r=0.5$.
+A distinct difference in distribution is visible between classification (a, b) and regression (c, d) datasets.
+Observations and Interpretation: The histograms in Figure 7 reveal a stark contrast between task types.
+In the node classification datasets (Cora, Citeseer), a significant number of nodes retain their 2nd-hop neighborhood entirely (lost fraction close to 0).
+Conversely, in the node regression datasets (Squirrel, Chameleon), the vast majority of nodes lose a highly significant portion of their 2nd-hop neighborhood (lost fraction close to 1).
+Given the superior performance of FIT-GNN on these regression tasks, we interpret this structural loss as a form of implicit adversarial pruning.
+In specific heterophilic graphs, long-range information (such as 2nd-hop neighbors) acts as noise or an adversarial signal.
+The coarsening algorithms utilized by FIT-GNN implicitly filter out this distant noise.
+This pruning allows the model to fully exploit the low-variance, highly homogeneous local structures identified in Section G, thereby drastically reducing regression error.